### v1

# Large Stepsize Gradient Descent for Non-Homogeneous Two-Layer Networks: Margin Improvement and Fast Optimization

Abstract
The typical training of neural networks using large stepsize gradient descent (GD) under the logistic loss often involves two distinct phases, where the empirical risk oscillates in the first phase but decreases monotonically in the second phase. We investigate this phenomenon in two-layer networks that satisfy a near-homogeneity condition. We show that the second phase begins once the empirical risk falls below a certain threshold, dependent on the stepsize. Additionally, we show that the normalized margin grows nearly monotonically in the second phase, demonstrating an implicit bias of GD in training non-homogeneous predictors. If the dataset is linearly separable and the derivative of the activation function is bounded away from zero, we show that the average empirical risk decreases, implying that the first phase must stop in finite steps. Finally, we demonstrate that by choosing a suitably large stepsize, GD that undergoes this phase transition is more efficient than GD that monotonically decreases the risk. Our analysis applies to networks of any width, beyond the well-known neural tangent kernel and mean-field regimes.
1 Introduction
Neural networks are mostly optimized by gradient descent (GD) or its variants. Understanding the behavior of GD is one of the key challenges in deep learning theory. However, there is a nonnegligible discrepancy between the GD setups in theory and in practice. In theory, GD is mostly analyzed with relatively small stepsizes such that its dynamics are close to the continuous gradient flow dynamics, although a few exceptions will be discussed later. However, in practice, GD is often used with a relatively large stepsize, with behaviors significantly deviating from that of small stepsize GD or gradient flow. Specifically, notice that small stepsize GD (hence also gradient flow) induces monotonically decreasing empirical risk, but in practice, good optimization and generalization performance is usually achieved when the stepsize is large and the empirical risk oscillates [see Wu and Ma, 2018, Cohen et al., 2020, for example]. Therefore, it is unclear which of the theoretical insights drawn from analyzing small stepsize GD apply to large stepsize GD used practically.
The behavior of small stepsize GD is relatively well understood. For instance, classical optimization theory suggests that GD minimizes convex and $L$-smooth functions if the stepsize $\tilde{\eta}$ is well below $2/L$, with a convergence rate of $\mathcal{O}(1/(\tilde{\eta}t))$, where $t$ is the number of steps [Nesterov, 2018]. More recently, Soudry et al. [2018], Ji and Telgarsky [2018] show an implicit bias of small stepsize GD in logistic regression with separable data, where the direction of the GD iterates converges to the max-margin direction. Subsequent works extend their implicit bias theory from linear model to homogenous networks [Lyu and Li, 2020, Chizat and Bach, 2020, Ji and Telgarsky, 2020]. These theoretical results all assume the stepsize of GD is small (and even infinitesimal) such that the empirical risk decreases monotonically and, therefore cannot be directly applied to large stepsize GD used in practice.
More recently, large stepsize GD that induces oscillatory risk has been analyzed in simplified setups [see Ahn et al., 2023, Zhu et al., 2022, Kreisler et al., 2023, Chen and Bruna, 2023, Wang et al., 2022a, Wu et al., 2023, 2024, for an incomplete list of references]. In particular, in logistic regression with linearly separable data, Wu et al. [2023] showed that the implicit bias of GD (that maximizes the margin) holds not only for small stepsizes [Soudry et al., 2018, Ji and Telgarsky, 2018] but also for an arbitrarily large stepsize. In the same problem, Wu et al. [2024] further showed that large stepsize GD that undergoes risk oscillation can achieve an $\tilde{\mathcal{O}}(1/t^{2})$ empirical risk, whereas small stepsize GD that monotonically decreases the empirical risk must suffer from a $\Omega(1/t)$ empirical risk. Nonetheless, these theories of large stepsize GD are limited to relatively simple setups such as linear models. The theory of large stepsize GD for non-linear networks is underdeveloped.
This work fills the gap by providing an analysis of large stepsize GD for non-linear networks. In the following, we set up our problem formally and summarize our contributions.
(a) Empirical risk
(b) Normalized margin
(c) Test accuracy
Figure 1: The behavior of (GD) for optimizing a non-homogenous MLP with tanh activation function on a subset of CIFAR-10 dataset. We randomly sample $1,979$ data with labels “airplane” and “automobile” from CIFAR-10 dataset. The normalized margin is defined by 3. The blue curves correspond to GD with a large stepsize $\tilde{\eta}=1$, where the empirical risk oscillates in the first phase but decreases monotonically in the second phase. The orange curves correspond to GD with a small stepsize $\tilde{\eta}=0.02$, where the empirical risk decreases monotonically. Furthermore, Figure 1(b) suggests the normalized margins of both two curves increase and converge in the stable phases. Finally, Figure 1(c) suggests that large stepsize achieves a better test accuracy, consistent with larger-scale learning experiment [Goyal et al., 2017]. More details can be found in Section 5.
Setup.
Consider a binary classification dataset $(\mathbf{x}_{i},y_{i})_{i=1}^{n}$, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ is a feature vector and $y_{i}\in\{\pm 1\}$ is a binary label. For simplicity, we assume $\|\mathbf{x}_{i}\|\leq 1$ for all $i$ throughout the paper. For a predictor $f$, the empirical risk under logistic loss is defined as
$L(\mathbf{w}):=\frac{1}{n}\sum_{i=1}^{n}\ell(y_{i}f(\mathbf{w};\mathbf{x}_{i})),\quad\ell(t):=\log(1+e^{-t}).$ (1)
Here, the predictor $f(\mathbf{w};\cdot):\mathbb{R}^{d}\mapsto\mathbb{R}$ is parameterized by trainable parameters $\mathbf{w}$ and is assumed to be continuously differentiable with respect to $\mathbf{w}$. The predictor is initialized from $\mathbf{w}_{0}$ and then trained by gradient descent (GD) with a constant stepsize $\tilde{\eta}>0$, that is,
$\mathbf{w}_{t+1}:=\mathbf{w}_{t}-\tilde{\eta}\nabla L(\mathbf{w}_{t}),\quad t\geq 0.$ (GD)
We are interested in a non-linear predictor $f$ and a large stepsize $\tilde{\eta}$. A notable example in our theory is two-layer networks with Lipschitz, smooth, and nearly homogenous activations (see 2). Note that minimizing $L(\mathbf{w})$ is a non-convex problem in general.
Observation.
Empirically, large stepsize GD often undergoes a phase transition, where the empirical risk defined in 1 oscillates in the first phase but decreases monotonically in the second phase (see empirical evidence in Appendix A in [Cohen et al., 2020] and a formal proof in [Wu et al., 2024] for linear predictors). This is illustrated in Figure 1 (the experimental setup is described in Section 5). We follow Wu et al. [2024] and call the two phases the edge of stability (EoS) phase [name coined by Cohen et al., 2020] and the stable phase, respectively.
Contributions.
We prove the following results for large stepsize GD for training non-linear predictors under logistic loss.
  1. 1.
    For Lipschitz and smooth predictor $f$ trained by GD with stepsize $\tilde{\eta}$, we show that as long as the empirical risk is below a threshold depending on $\tilde{\eta}$, GD monotonically decreases the empirical risk (see Theorem 2.2). This result extends the stable phase result in Wu et al. [2024] from linear predictors to non-linear predictors, demonstrating the generality of the existence of a stable phase.
  2. 2.
    Assuming that GD enters the stable phase, if in addition the preditor has a bounded homogenous error (see 1C), we show that the normalized margin induced by GD, $\min_{i}y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i})/\|\mathbf{w}_{t}\|$, nearly monotonically increases (see Theorem 2.2). To the best of our knowledge, this is the first characterization of implicit bias of GD for non-homogenous predictors. In particular, our theory covers two-layer networks with commonly used activations functions (which are often non-homogenous) that cannot be covered by existing results [Lyu and Li, 2020, Ji and Telgarsky, 2020, Chizat and Bach, 2020].
  3. 3.
    Under additional technical assumptions (the dataset is linearly separable and the derivative of the activation function is bounded away from zero), we show that the initial EoS phase must stop in $\mathcal{O}(\tilde{\eta})$ steps and GD transits to the stable phase afterwards. Furthermore, by choosing a suitably large stepsize, GD achieves a $\tilde{\mathcal{O}}(1/t^{2})$ empirical risk after $t$ steps. In comparison, GD that converges monotonically incurs an $\Omega(1/t)$ risk. This result indicates an optimization benefit of using large stepsize and generalizes the results in [Wu et al., 2024] from linear predictors to neural networks.
2 Stable Phase and Margin Improvement
In this section, we present our results for the stable phase of large stepsize GD in training non-linear predictors. Specifically, our results apply to non-linear predictors that are Lipschitz, smooth, and nearly homogeneous, as described by the following assumption.
Assumption 1 (Model conditions).
Consider a predictor $f(\mathbf{w};\mathbf{x}_{i})$, where $\mathbf{x}_{i}$ is one of the feature vectors in the training set.
  1. A.
    Lipschitzness. Assume there exists $\rho>0$ such that for every $\mathbf{w}$, $\sup_{i}\|\nabla_{\mathbf{w}}f(\mathbf{w};\mathbf{x}_{i})\|\leq\rho$.
  2. B.
    Smoothness. Assume there exists $\beta>0$ such that for all $\mathbf{w},\mathbf{v}$,
    $\|\nabla f(\mathbf{w};\mathbf{x}_{i})-\nabla f(\mathbf{v};\mathbf{x}_{i})\|\leq\beta\|\mathbf{w}-\mathbf{v}\|,\quad i=1,\dots,n.$
  3. C.
    Near homogeneity. Assume there exists $\kappa>0$ such that for every $\mathbf{w}$,
    $|f(\mathbf{w};\mathbf{x}_{i})-\langle\nabla_{\mathbf{w}}f(\mathbf{w};\mathbf{x}_{i}),\mathbf{w}\rangle|\leq\kappa,\quad i=1,\dots,n.$
1A and 1B are commonly used conditions in the optimization literature. If $\kappa=0$, then 1C requires the predictor to be exactly $1$-homogenous. Our 1C allows the predictor to have a bounded homogenous error. It is clear that linear predictors $f(\mathbf{w};\mathbf{x}):=\mathbf{w}^{\top}\mathbf{x}$ satisfy Assumption 1 with $\rho=\sup_{i}\|\mathbf{x}_{i}\|\leq 1$, $\beta=0$, and $\kappa=0$. Another notable example is two-layer networks given by
$f(\mathbf{w};\mathbf{x}):=\frac{1}{m}\sum_{j=1}^{m}a_{j}\phi(\mathbf{x}^{\top}\mathbf{w}^{(j)}),\quad\mathbf{w}^{(j)}\in\mathbb{R}^{d},\quad j=1,\dots,m,$ (2)
where we assume $a_{j}\in\{\pm 1\}$ are fixed and $\mathbf{w}=(\mathbf{w}^{(j)})_{j=1}^{m}\in\mathbb{R}^{md}$ are the trainable parameters. We define two-layer networks with the mean-field scaling [Song et al., 2018, Chizat and Bach, 2020]. However, our results hold for any width. The effect of rescaling the model will be discussed in Section 4. The following example shows that Assumption 1 covers two-layer networks with many commonly used activations $\phi(\cdot)$. The proof is provided in Section E.1.
Example 2.1 (Two-layer networks).
Two-layer networks defined in (2) with the following activation functions satisfy Assumption 1 with the described constants:
  • GELU. $\phi(x):=\frac{x}{2}\mathrm{erf}\big{(}1+(x/\sqrt{2})\big{)}$ with $\kappa={e^{-1/2}}/{\sqrt{2\pi}}$, $\beta=2/m$, and $\rho=(\sqrt{2\pi}+e^{-1/2})/\sqrt{2\pi m}$.
  • Softplus. $\phi(x):=\log(1+e^{x})$ with $\kappa=\log{2}$, $\beta=1/m$, and $\rho=1/\sqrt{m}$.
  • Sigmoid. $\phi(x):={1}/{(1+e^{-x})}$ with $\kappa=1$, $\beta=2/m$, and $\rho=1/\sqrt{m}$.
  • Tanh. $\phi(x):=\tanh(x)$ with $\kappa=5$, $\beta=2/m$, and $\rho=1/\sqrt{m}$.
  • SiLU. $\phi(x):={x}/{(1+e^{-x})}$ with $\kappa=1$, $\beta=4/m$, and $\rho=2/\sqrt{m}$.
  • Huberized ReLU [Chatterji et al., 2021]. For a fixed $h>0$,
    $\phi(x)\coloneqq\begin{cases}0&x<0,\\ \frac{x^{2}}{2h}&0\leq x\leq h,\\ x-\frac{h}{2}&x>h,\end{cases}$
    with $\kappa=h/2$, $\beta=1/(hm)$, and $\rho=1/\sqrt{m}$.
Margin for nearly homogenous predictors.
For a nearly homogenous predictor $f(\mathbf{w};\cdot)$ (see 1C), we define its normalized margin (or margin for simplicity) as
$\bar{\gamma}(\mathbf{w}):=\frac{\min_{i\in[n]}y_{i}f(\mathbf{w};\mathbf{x}_{i})}{\|\mathbf{w}\|}.$ (3)
A large normalized margin $\bar{\gamma}(\mathbf{w})$ guarantees the prediction of each sample is away from the decision boundary. The normalized margin 3 is introduced by Lyu and Li [2020] for homogenous predictors. However, we show that the same notion is also well-defined for non-homogenous predictors that satisfy 1C. The next theorem gives sufficient conditions for large stepsize GD to enter the stable phase in training non-homogenous predictors and characterizes the increase of the normalized margin. The proof of Theorem 2.2 is deferred to Appendix A.
Theorem 2.2 (Stable phase and margin improvement).
Consider GD with stepsize $\tilde{\eta}$ on a predictor $f(\mathbf{w};\mathbf{x})$ that satisfies 1A and 1B. If there exists $r\geq 0$ such that
$L(\mathbf{w}_{r})\leq\frac{1}{\tilde{\eta}(2\rho^{2}+\beta)},$ (4)
then GD is in the stable phase for $t\geq r$, that is, $(L(\mathbf{w}_{t}))_{t\geq r}$ decreases monotonically. If additionally the predictor satisfies 1C and there exists $s\geq 0$ such that
$L(\mathbf{w}_{s})\leq\min\Big{\{}\frac{1}{e^{\kappa+2}2n},\ \frac{1}{\tilde{\eta}(4\rho^{2}+2\beta)}\Big{\}},$ (5)
we have the following for $t\geq s$:
  • Risk convergence. $L(\mathbf{w}_{t})=\Theta(1/t)$, where the constants depend on $L(\mathbf{w}_{s}),\|\mathbf{w}_{s}\|,\tilde{\eta}$ and $\rho$.
  • Parameter increase. $\|\mathbf{w}_{t+1}\|\geq\|\mathbf{w}_{t}\|$ and $\|\mathbf{w}_{t}\|=\Theta(\log(t))$, where the constants depend on $n,\rho,\kappa,\tilde{\eta},\|\mathbf{w}_{s}\|,L(\mathbf{w}_{s})$ and $f(\mathbf{w}_{s};\mathbf{x}_{i})$.
  • Margin improvement. There exists a modified margin function $\gamma^{c}(\mathbf{w})$ such that
    • $\gamma^{c}(\mathbf{w}_{t})$ is increasing and bounded.
    • $\gamma^{c}(\mathbf{w}_{t})$ is a multiplicative approximiator of $\bar{\gamma}(\mathbf{w}_{t})$, that is, there exists $c>0$ such that
      $\gamma^{c}(\mathbf{w}_{t})\leq\bar{\gamma}(\mathbf{w}_{t})\leq\bigg{(}1+\frac{c}{\log(1/{L(\mathbf{w}_{t}))}}\bigg{)}\gamma^{c}(\mathbf{w}_{t}),\quad t\geq s.$
    As a direct consequence, $\lim_{t\to\infty}\bar{\gamma}(\mathbf{w}_{t})=\lim_{t\to\infty}\gamma^{c}(\mathbf{w}_{t})$.
Theorem 2.2 shows that for an arbitrarily large stepsize $\tilde{\eta}$, GD must enter the stable phase if the empirical risk falls below a threshold depending on $\tilde{\eta}$ given by 4. In the next section, We will show that under additional technical conditions, large stepsize GD is guaranteed to satisfy 4 even though with an initially oscillatory risk.
Furthermore, for nearly homogenous predictors, Theorem 2.2 shows that under a stronger risk threshold condition 5, the risk must converge at a $\Theta(1/t)$ rate and that the normalized margin nearly monotonically increases. This demonstrates an implicit bias of GD, even when used with a large stepsize and the trained predictor is non-homogenous.
Our Theorem 2.2 makes several important extensions compared to existing results. First, Theorem 2.2 suggests that the stable phase happens for general non-linear predictors such as two-layer networks, while the work by Wu et al. [2024] only studied the stable phase for linear predictors. Second, the margin improvement is only known for small (and even infinitesimal) stepsize GD and homogenous predictors [Lyu and Li, 2020, Chizat and Bach, 2020, Ji and Telgarsky, 2020]. To the best of our knowledge, Theorem 2.2 is the first implicit bias result that covers large stepsize GD and non-homogenous predictors.
From a technical perspective, our proof uses techniques introduced by Lyu and Li [2020] for analyzing homogenous predictors. Our main innovations are constructing new auxiliary margin functions that can deal with errors caused by large stepsize and non-homogeneity. More details are discussed in Section A.3.
3 Edge of Stability Phase
Our stable phase results in Theorem 2.2 require the risk to be below a certain threshold (see 4 and 5). In this section, we show that the risk can indeed be below the required threshold, even when GD is used with large stepsize. Recall that minimizing the empirical risk with a non-linear predictor is non-convex, therefore solving it by GD is hard in general. We make additional technical assumptions to conquer the challenges caused by non-convexity. We conjecture that these technical assumptions are not necessary and can be relaxed.
We focus on two-layer networks 2. We make the following assumptions on the activation function.
Assumption 2 (Activation function conditions).
In the two-layer network 2, let the activation function $\phi:\mathbb{R}\to\mathbb{R}$ be continuously differentiable. Moreover,
  1. A.
    Derivative condition. Assume there exists $10<\alpha<$ 1 such that $\alpha\leq|\phi^{\prime}(z)|\leq 1$.
  2. B.
    Smoothness. Assume there exists $\tilde{\beta}>0$ such that for all $x,y\in\mathbb{R}$, $|\phi^{\prime}(x)-\phi^{\prime}(y)|\leq\tilde{\beta}|x-y|.$
  3. C.
    Near homogeneity. Assume there exists $\kappa>0$ such that for every $z\in\mathbb{R}$, $|\phi(z)-\phi^{\prime}(z)z|\leq\kappa$.
Recall that $\sup_{i}\|\mathbf{x}_{i}\|\leq 1$. One can then check by direct computation that, under Assumption 2, two-layer networks 2 satisfy Assumption 1 with $\rho=1/\sqrt{m}$, $\beta=\tilde{\beta}/m$, and $\kappa=\kappa$.
2B and 2C cover many commonly used activation functions. In 2A, we assume $|\phi^{\prime}(z)|\leq 1$. This is just for the simplicity of presentation and our results can be easily generalized to allow $|\phi^{\prime}(z)|\leq C$ for a constant $C>0$. The other condition in 2A, $|\phi^{\prime}(z)|\geq\alpha$, however, is non-trivial. This condition is widely used in literature [see Brutzkus et al., 2018, Frei et al., 2021, and references thereafter] to facilitate GD analysis. Technically, this condition guarantees that each neuron in the two-layer network 2 will always receive a non-trivial gradient in the GD update; otherwise, neurons may be frozen during the GD update. Furthermore, commonly used activation functions can be combined with an identity map to satisfy 2A. This is formalized in the following example. The proof is provided in Section E.2.
Example 3.1 (Leaky activation functions).
Fix $10.5\leq c<$ 1. For each $\phi$ in Example 2.1 except the Huberized ReLU, its modification $\tilde{\phi}(x):=cx+(1-c)/4\cdot\phi(x)$ satisfies Assumption 2 with $\kappa=1$, $\alpha=0.25$, and $\tilde{\beta}=1$. The modification of the Huberized ReLU satisfies Assumption 2 with $\kappa=h/2$, $\alpha=0.5$, and $\tilde{\beta}=1/4h$. In particular, the modification of softplus can be viewed as a smoothed leaky ReLU.
For the technical difficulty of non-convex optimization, we also need to assume a linearly separable dataset to conduct our EoS phase analysis.
Assumption 3 (Linear separability).
Assume there is a margin $\gamma>0$ and a unit vector $\mathbf{w}_{*}$ such that $y_{i}\mathbf{x}_{i}^{\top}\mathbf{w}_{*}\geq\gamma$ for every $i=1,\ldots,n$.
The following theorem shows that when GD is used with large stepsizes, the average risk must decrease even though the risk may oscillate locally. The proof of Theorem 3.2 is defered to Section 3.
Theorem 3.2 (The EoS phase for two-layer networks).
Under Assumption 3, consider GD on two-layer networks 2 that satisfy 2A and 2C. Denote the stepsize by $\tilde{\eta}:=m\eta$, where $m$ is the network width and $\eta$ can be arbitrarily large. Then for every $t>0$, we have
$\frac{1}{t}\sum_{k=0}^{t-1}L(\mathbf{w}_{k})\leq\frac{1+8\log^{2}(\gamma^{2}\eta t)/\alpha^{2}+8\kappa^{2}/\alpha^{2}+\eta^{2}}{\gamma^{2}\eta t}+\frac{\|\mathbf{w}_{0}\|^{2}}{m\eta t}=\mathcal{O}\bigg{(}\frac{\log^{2}(\eta t)+\eta^{2}}{\eta t}\bigg{)}.$
Theorem 3.2 suggests that the average risk of training two-layer networks decreases even when GD is used with large stepsize. Consequently, the risk thresholds 4 and 5 for GD to enter the stable phase must be satisfied after a finite number of steps. This will be discussed in depth in the next section.
Compared to Theorem 1 in [Wu et al., 2024], Theorem 3.2 extends their EoS phase bound from linear predictors to two-layer networks.
4 Phase Transition and Fast Optimization
For two-layer networks trained by large stepsize GD, Theorem 3.2 shows that the average risk must decrease over time. Combining this with Theorem 2.2, GD must enter the stable phase in finite steps, and the loss must converge while the normalized margin must improve.
However, a direct application of Theorem 3.2 only leads to a suboptimal bound on the phase transition time. Motivated by Wu et al. [2024], we establish the following sharp bound on the phase transition time by tracking the gradient potential (see Lemma C.3). The proof of Theorem 4.1 is deferred to Appendix C.
Theorem 4.1 (Phase transition and stable phase for two-layer networks).
Under Assumption 3, consider GD on two-layer networks 2 that satisfy Assumption 2. Clearly, the two-layer networks also satisfy Assumption 1 with $\rho=1/\sqrt{m}$, $\beta=\tilde{\beta}/m$, and $\kappa=\kappa$. Denote the stepsize by $\tilde{\eta}:=m\eta$, where $m$ is the network width and $\eta>0$ can be arbitrarily large.
  • Phase transition time. There exists $s\leq\tau$ such that (5) in Theorem 2.2 holds, where
    $\tau:=\frac{128(1+4\kappa)}{\alpha^{2}}\max\bigg{\{}c_{1}\eta,c_{2}n,e,\frac{c_{2}\eta+c_{1}n}{\eta}\log\frac{c_{2}\eta+c_{1}n}{\eta},\frac{(c_{2}\eta+c_{1}n)}{\eta}\cdot\frac{\|\mathbf{w}_{0}\|}{\sqrt{m}}\bigg{\}},$
    where $c_{1}:=4e^{\kappa+2}$ and $c_{2}:=(8+4\tilde{\beta})$. Therefore GD is in the stable phase from $s$ onwards.
  • Explicit risk bound in the stable phase. We have $(L(\mathbf{w}_{t}))_{t\geq s}$ monotonically decreases and
    $L(\mathbf{w}_{t})\leq\frac{2}{\alpha^{2}\gamma^{2}\eta(t-s)},\quad t\geq s.$
Theorems 2.2, 3.2 and 4.1 together characterize the behaviors of large stepsize GD in training two-layer networks. Specifically, large stepsize GD may induce an oscillatory risk in the beginning; but the averaged empirical risk must decrease (Theorem 3.2). After the empirical risk falls below a certain stepsize-dependent threshold, GD enters the stable phase, where the risk decreases monotonically (Theorem 4.1). Finally, the normalized margin 3 induced by GD increases nearly monotonically as GD stays in the stable phase (Theorem 2.2).
Fast optimization.
Our bounds for two-layer networks are comparable to those for linear predictors shown by Wu et al. [2024]. Specifically, when used with a larger stepsize, GD achieves a faster optimization in the stable phase but stays longer in the EoS phase. Choosing a suitably large stepsize that balances the steps in EoS and stable phases, we obtain an accelerated empirical risk in the following corollary. The proof is included in Section C.2.
Corollary 4.2 (Acceleration of large stepsize).
Under the same setup as in Theorem 4.1, consider (GD) with a given budget of $T$ steps such that
[…]
where $c_{1}:=4e^{\kappa+2}$ and $c_{2}:=(8+4\tilde{\beta})$ are as in Theorem 4.1. Then for stepsize $\tilde{\eta}:=\eta m$, where
$\eta:=\frac{\alpha^{2}\gamma^{2}}{256(1+4\kappa)c_{2}}T,$
we have $\tau\leq T/2$ and
$L(\mathbf{w}_{T})\leq\frac{2048(1+4\kappa)c_{2}}{\alpha^{4}\gamma^{4}}\cdot\frac{1}{T^{2}}=\mathcal{O}\bigg{(}\frac{1}{T^{2}}\bigg{)}.$
Theorem 4.1 and Corollary 4.2 extend Theorem 1 and Corollary 2 in Wu et al. [2024] from linear predictors to two-layer networks. Another notable difference is that we obtain a sharper stable phase bound (and thus a better acceleration bound) compared to theirs, where we remove a logarithmic factor through a more careful analysis.
Corollary 4.2 suggests an accelerated risk bound of $\mathcal{O}(1/T^{2})$ by choosing a large stepsize that balances EoS and stable phases. We also show the following lower bound, showing that such acceleration is impossible if GD does not enter the EoS phase. The proof is included in Section C.3.
Theorem 4.3 (Lower bound in the classical regime).
Consider (GD) with initialization $\mathbf{w}_{0}=0$ and stepsize $\tilde{\eta}>0$ for a two-layer network 2 satisfying Assumption 2. Suppose the training set is given by
$\mathbf{x}_{1}=(\gamma,\sqrt{1-\gamma^{2}}),\quad\mathbf{x}_{2}=(\gamma,-\sqrt{1-\gamma^{2}}/2),\quad y_{1}=y_{2}=1,\quad 0<\gamma<0.1.$
It is clear that $\left(\mathbf{x}_{i},y_{i}\right)_{i=1,2}$ satisfy Assumption 3. If $\left(L\left(\mathbf{w}_{t}\right)\right)_{t\geq 0}$ is non-increasing, then
$L\left(\mathbf{w}_{t}\right)\geq c_{0}/t,\quad t\geq 1$
where $c_{0}>0$ is a function of $(\alpha,\phi,\mathbf{x}_{1},\mathbf{x}_{2},\gamma,\kappa,\beta)$ but is independent of $t$ and $\tilde{\eta}$.
Effect of model rescaling.
We conclude this section by discussing the impact of rescaling the model. Specifically, we replace the two-layer network in the mean-field scaling 2 by the following
$f(\mathbf{w};\mathbf{x}):=b\cdot\frac{1}{m}\sum_{j=1}^{m}a_{j}\phi(\mathbf{x}^{\top}\mathbf{w}^{(j)}),$
and evaluate the impact of the scaling factor $b$ on our results. By choosing the optimal stepsize that balances the EoS and stable phases as in Corollary 4.2, we optimize the risk bound obtained by GD with a fixed budget of $T$ steps and get the following bound. Detailed derivations are deferred to Appendix D.
$L(\mathbf{w}_{T})=\begin{cases}\mathcal{O}(1/T^{2})&\text{if }b\geq 1,\\ \mathcal{O}(b^{-3}/T^{2})&\text{if }b<1.\end{cases}$
This suggests that as long as $b\geq 1$, we get the same acceleration effect. In particular, the mean-field scaling $b=1$ [Song et al., 2018, Chizat and Bach, 2020] and the neural tangent kernel (NTK) scaling $b=\sqrt{m}$ [Du et al., 2018, Jacot et al., 2018] give the same $\mathcal{O}(1/T^{2})$ acceleration effect. An NTK analysis of large stepsize is included in [Wu et al., 2024] and their conclusion is consistent with ours. Finally, we remark that our analysis holds for any width $m$ and uses techniques different from the mean-field or NTK methods. However, our acceleration analysis only allows linearly separable datasets.
5 Experiments
We conduct three sets of experiments to validate our theoretical insights. In the first set, we use a subset of the CIFAR-10 dataset [Krizhevsky et al., 2009], which includes 1,979 randomly selected samples from the ”airplane” and ”automobile” classes. Our model is a multilayer perceptron (MLP) with two trainable layers and tanh activation functions, with a hidden dimension of 100. The MLP is trained using gradient descent with random initialization, as described in GD. The results are shown in Figures 1(a), 1(c) and 1(b).
In the second set of experiments, we consider an XOR dataset consisting of four samples:
$\mathbf{x}_{1}=(-1,-1),y_{1}=1;\ \mathbf{x}_{2}=(1,1),y_{2}=1;\ \mathbf{x}_{3}=(1,-1),y_{3}=-1;\ \mathbf{x}_{4}=(-1,1),y_{4}=-1.$
The above XOR dataset is not linearly separable. We test GD with different stepsizes on a two-layer network 2 with the leaky softplus activation (see Example 3.1 with $c=0.5$). The network width is $m=20$. The initialization is random. The results are presented in Figures 2(a), 2(b) and 2(c).
In the third set of experiments, we consider the same task as in the first set of experiments, but we test GD with different stepsizes on a two-layer network 2 with the softplus activation. The network width is $m=40$. The initialization is random. The results are presented in Figures 2(d), 2(e) and 2(f).
Margin improvement.
Figures 1(b), 2(c) and 2(f) show that the normalized margin nearly monotonically increases once gradient descent (GD) enters the stable phase, regardless of step size. This observation aligns with our theoretical findings in Theorem 2.2.
Fast optimization.
From Figures 1(a), 2(a) and 2(d), we observe that after GD enters the stable phase, a larger stepsize consistently leads to a smaller empirical risk compared to the smaller stepsizes, which is consistent with our Theorem 4.1 and Corollary 4.2. Besides, Figures 2(b) and 2(e) suggest that, asymptotically, GD converges at a rate of $\mathcal{O}(1/(\tilde{\eta}t))=\mathcal{O}(1/(\eta t))$ (The width of networks is fixed), which verifies the sharpness of our stable phase bound in Theorem 4.1.
(a) Empirical risk, XOR.
(b) Asymptotic rate, XOR.
(c) Normalized margin, XOR.
(d) Empirical risk, CIFAR-10.
(e) Asymptotic rate, CIFAR-10.
(f) Normalized margin, CIFAR-10.
Figure 2: Behavior of (GD) for two-layer networks 2 with leaky softplus activation function (see Example 3.1 with $c=0.5$). We consider an XOR dataset and a subset of CIFAR-10 dataset. In both cases, we observe that (1) GD with a large stepsize achieves a faster optimization compared to GD with a small stepsize, (2) the asymptotic convergence rate of the empirical risk is $\mathcal{O}(1/(\tilde{\eta}t))$, and (3) in the stable phase, the normalized margin (nearly) monotonically increases. These observations are consistent with our theoretical understanding of large stepsize GD. More details about the experiments are explained in Section 5.
6 Related Works
In this section, we discuss related papers.
Small stepsize and implicit bias.
For logistic regression on linearly separable data, Soudry et al. [2018], Ji and Telgarsky [2018] showed that the direction of small stepsize GD converges to the max-margin solution. Their results were later extended by Gunasekar et al. [2017, 2018], Nacson et al. [2019c, a, b], Ji et al. [2021], Lyu and Li [2020], Ji and Telgarsky [2020], Chizat and Bach [2020] to other algorithms and non-linear models. However, in all of their analysis, the stepsize of GD needs to be small such that the empirical risk decreases monotonically. In contrast, our focus is GD with a large stepsize that induces non-monotonic risk.
Large stepsize and EoS.
In practice, large stepsizes are often preferred when using GD to train neural networks to achieve effective optimization and generalization performance [see Wu and Ma, 2018, Cohen et al., 2020, and references therein]. In such scenarios, the empirical risk often oscillates in the beginning. This phenomenon is named edge of stability (EoS) by Cohen et al. [2020]. The theory of EoS is mainly studied in relatively simplified cases such as one- or two-dimensional functions [Zhu et al., 2022, Chen and Bruna, 2023, Ahn et al., 2022, Kreisler et al., 2023, Wang et al., 2023], linear model [Wu et al., 2023, 2024], matrix factorization [Wang et al., 2022a, Chen and Bruna, 2023], scale-invariant networks [Lyu et al., 2022], for an incomplete list of references. Compared to them, we focus on a more practical setup of training two-layer non-linear networks with large stepsize GD. There are some general theories of EoS subject to subtle assumptions [Kong and Tao, 2020, Ahn et al., 2022, Ma et al., 2022, Damian et al., 2022, Wang et al., 2022b], which are not directly comparable to ours.
In what follows, we make a detailed discussion about papers that directly motivate our work [Lyu and Li, 2020, Ji and Telgarsky, 2020, Chatterji et al., 2021, Wu et al., 2023, 2024].
Comparison with Lyu and Li [2020], Ji and Telgarsky [2020].
Both results in [Lyu and Li, 2020, Ji and Telgarsky, 2020] focused on $L$-homogenous networks. Specifically, Lyu and Li [2020] showed that a modified version of normalized margin (see 3) induced by GD with small stepsize (such that the risk decreases monotonically) increases, with limiting points of $\{\mathbf{w}_{t}/\|\mathbf{w}_{t}\|\}_{t=1}^{\infty}$ converging to KKT points of a margin-maximization problem. Under additional o-minimal conditions, Ji and Telgarsky [2020] showed that gradient flow converges in direction. Our work is different from theirs in two aspects. First, we allow GD with a large stepsize that may cause risk oscillation. Second, our theory covers non-homogenous predictors, which include two-layer networks with many commonly used activation functions beyond the scope of [Lyu and Li, 2020, Ji and Telgarsky, 2020]. Compared to Lyu and Li [2020], Ji and Telgarsky [2020], we only show the improvement of the margin, and our theory is limited to nearly $1$-homogenous predictors (2C). It remains open to show directional convergence and to extend our near $1$-homogenity condition to a “near $L$-homogeneity” condition.
Comparison with Chatterji et al. [2021].
The work by Chatterji et al. [2021] studies the convergence of GD in training deep networks under logistic loss. Their results are related to ours as we both consider networks with nearly homogeneous activations and we both have a stable phase analysis (although this is not explicitly mentioned in their paper). However, our results are significantly different from theirs. Specifically, in our notation, they require the homogenous error $\kappa$ (see 2C) to be smaller than $\mathcal{O}(\log(1/L(\mathbf{w}_{s}))/\|\mathbf{w}_{s}\|)\approx\mathcal{O}(\bar{\gamma}(\mathbf{w}_{s}))$, where $s$ is the time for GD to enter the stable phase. Note that the margin when GD enters the stable phase could be arbitrarily small. In comparison, we only require the homogenous error to be bounded by a constant. As a consequence, we can handle many commonly used activation functions (see Example 2.1) while they can only handle the Huberized ReLU with a small $h$ in Example 2.1. Moreover, they require the stepsize $\tilde{\eta}$ to be smaller than $\mathcal{O}(\kappa/\|\mathbf{w}_{s}\|^{8})$, thus they only allow very small stepsize. In contrast, we allow $\tilde{\eta}$ to be an arbitrarily large constant.
Comparison with Wu et al. [2023, 2024].
The works by Wu et al. [2023, 2024] directly motivate our paper. In particular, for logistic regression on linearly separable data, Wu et al. [2023] showed margin maximization of GD with large stepsize and Wu et al. [2024] showed fast optimization of GD with large stepsize. Our work can be viewed as an extension of [Wu et al., 2023, 2024] from linear predictors to non-linear predictors such as two-layer networks. Besides, our results for margin improvement and convergence within the stable phase (Theorem 2.2) hold for the general dataset, while their results strongly rely on the linear separability of the dataset.
7 Conclusion
We provide a theory of large stepsize gradient descent (GD) for training non-homogeneous predictors such as two-layer networks using the logistic loss function. Our analysis explains the empirical observations: large stepsize GD often reveals two distinct phases in the training process, where the empirical risk oscillates in the beginning but decreases monotonically subsequently. We show that the phase transition happens because the average empirical risk decreases despite the risk oscillation. In addition, we show that large stepsize GD improves the normalized margin in the long run, which extends the existing implicit bias theory for homogenous predictors to non-homogenous predictors. Finally, we show that large stepsize GD, by entering the initial oscillatory phase, achieves acceleration when minimizing the empirical risk.
Acknowledgements
We thank Fabian Pedregosa for his suggestions on an early draft. We gratefully acknowledge the support of the NSF for FODSI through grant DMS-2023505, of the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and #814639, and of the ONR through MURI award N000142112431.
Appendix A Stable Phase Analysis
In this section, we will prove results for a general smooth predictor $f(\mathbf{w};\mathbf{x})$ under the logistic loss in the stable phase. Before the proof, we introduce some notations here.
Notation.
We use the following notation to simplify the presentation.
  • $q_{i}(t)\coloneqq y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i}),\;q_{\min}(t)\coloneqq\min_{i\in[n]}q_{i}(t).$
  • $L_{t}\coloneqq L(\mathbf{w}_{t}),\;\rho_{t}:=\|\mathbf{w}_{t}\|_{2}.$
Then, we have the following expression:
$L(\mathbf{w}_{t})=\frac{1}{n}\sum_{i=1}^{n}\ell(q_{i}(t)).$
Here, we give a summary of this section. The proofs are organized into 5 parts.
  • In Section A.1, we characterize the decrease of loss $L_{t}$.
  • In Section A.2, we characterize the change of the parameter norm $\rho_{t}$.
  • In Section A.3, we show the convergence of the normalized margin $\bar{\gamma}(\mathbf{w}_{t})$.
  • In Section A.4, we characterize the sharp rates of loss $L_{t}$ and parameter norm $\rho_{t}$.
  • In Section A.5, we give the proof of Theorem 2.2.
A.1 Decrease of the Loss
In this section, we will show that the loss $L_{t}$ decreases monotonically in the stable phase. To begin with, we introduce the following definition which is another characterization of $\beta$-smoothness.
Definition 1 (Linearization error).
Given a continuously differentiable function $f:\mathbb{R}^{d}\to\mathbb{R}$ and two points $\mathbf{w},\mathbf{v}\in\mathbb{R}^{d}$, the linearization error of $f(\mathbf{v})$ with respect to $\mathbf{w}$ is:
$\xi[f](\mathbf{w},\mathbf{v}):=f(\mathbf{v})-f(\mathbf{w})-\nabla f(\mathbf{w})^{\top}(\mathbf{v}-\mathbf{w}).$
For a $\beta$-smooth function, standard convex optimization theory gives the following linearization error bound.
Fact A.1 (Linearization error of $\beta$-smooth function).
For a $\beta$-smooth function $f:\mathbb{R}^{d}\to\mathbb{R}$, we have
$\xi[f](\mathbf{w},\mathbf{v}):=f(\mathbf{v})-f(\mathbf{w})-\nabla f(\mathbf{w})^{\top}(\mathbf{v}-\mathbf{w})\leq\frac{\beta}{2}\|\mathbf{v}-\mathbf{w}\|_{2}^{2},\quad\text{for every $\mathbf{w}$ and $\mathbf{v}$}.$
We first show a stable phase bound for general smooth and Lipschitz predictors. The following is an extension of Lemma 10 in [Wu et al., 2024]. Since we do not require $f$ to be twice differentiable, extra efforts are needed.
Lemma A.2 (Self-boundedness of logistic loss).
For the logistic loss $\ell(z):=\log(1+\exp(-z))$, we have
$0\leq\ell(z)-\ell(x)-\ell^{\prime}(x)(z-x)\leq 2\ell(x)(z-x)^{2}$
for $|z-x|<1$.
Proof of Lemma A.2.
See the proof of Proposition 5 in [Wu et al., 2024]. The lower bound is by the convexity of $\ell(\cdot)$. ∎
The next lemma controls the decrease of the risk $L_{t}$.
Lemma A.3 (Decrease of $L_{t}$).
Suppose 1A and 1B hold. If $L(\mathbf{w}_{t})\leq\frac{1}{\tilde{\eta}\rho^{2}}$, then we have
$-\tilde{\eta}(1+\beta\tilde{\eta}L(\mathbf{w}_{t}))\|\nabla L(\mathbf{w}_{t})\|^{2}\leq L(\mathbf{w}_{t+1})-L(\mathbf{w}_{t})\leq-\tilde{\eta}(1-(2\rho^{2}+\beta)\tilde{\eta}L(\mathbf{w}_{t}))\|\nabla L(\mathbf{w}_{t})\|^{2}.$
Particularly, this indicates that if $L(\mathbf{w}_{t})\leq\frac{1}{\tilde{\eta}(2\rho^{2}+\beta)}$, then $L(\mathbf{w}_{t+1})\leq L(\mathbf{w}_{t})$.
Proof of Lemma A.3.
By 1A and 1B, we have $\|\nabla f\|_{2}\leq\rho$ and $f(\mathbf{w};\mathbf{x})$ is $\beta$-smooth as a function of $\mathbf{w}$. Therefore, for every $i\in[n]$ we have
$\left|q_{i}(t+1)-q_{i}(t)\right|$ $=\left|y_{i}(f\left(\mathbf{w}_{t+1};\mathbf{x}_{i}\right)-f\left(\mathbf{w}_{t};\mathbf{x}_{i}\right))\right|$
$=\left|\nabla f\left(\mathbf{w}_{t}+\theta\left(\mathbf{w}_{t+1}-\mathbf{w}_{t}\right);\mathbf{x}_{i}\right)^{\top}\left(\mathbf{w}_{t+1}-\mathbf{w}_{t}\right)\right|$ by intermediate value theorem
$\leq\rho\left\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\right\|$
$\leq\rho\tilde{\eta}\|\nabla L_{t}\|$ since $\mathbf{w}_{t+1}=\mathbf{w}_{t}-\tilde{\eta}\nabla L_{t}$
$\leq\rho^{2}\tilde{\eta}L_{t}\leq 1.$ since $\|\nabla L_{t}\|\leq L_{t}\rho$
Then by Lemma A.2, we have
$\ell(q_{i}(t+1))$ $\leq\ell(q_{i}(t))+\ell^{\prime}(q_{i}(t))(q_{i}(t+1)-q_{i}(t))+2\ell(q_{i}(t))(q_{i}(t+1)-q_{i}(t))^{2}$
$\leq\ell(q_{i}(t))+\ell^{\prime}(q_{i}(t))\langle y_{i}\nabla f(\mathbf{w}_{t};\mathbf{x}_{i}),\mathbf{w}_{t+1}-\mathbf{w}_{t}\rangle+|\ell^{\prime}(q_{i}(t))|\cdot|\xi[f](\mathbf{w}_{t},\mathbf{w}_{t+1})|$
$\quad+2\ell(q_{i}(t))(q_{i}(t+1)-q_{i}(t))^{2}$
since $q_{i}(t+1)-q_{i}(t)=\langle y_{i}\nabla f(\mathbf{w}_{t};\mathbf{x}_{i}),\mathbf{w}_{t+1}-\mathbf{w}_{t}\rangle+y_{i}\xi[f](\mathbf{w}_{t},\mathbf{w}_{t+1})$
$\leq\ell(q_{i}(t))+\ell^{\prime}(q_{i}(t))\langle y_{i}\nabla f(\mathbf{w}_{t};\mathbf{x}_{i}),\mathbf{w}_{t+1}-\mathbf{w}_{t}\rangle+\ell(q_{i}(t))(\beta+2\rho^{2})\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\|^{2}.$
by A.1 and the previous inequality
Taking an average over all data points, we have
$L_{t+1}\leq L_{t}-\tilde{\eta}\|\nabla L_{t}\|^{2}+(2\rho^{2}+\beta)\tilde{\eta}^{2}L_{t}\|\nabla L_{t}\|^{2},$
which is equivalent to
$L_{t+1}-L_{t}\leq-\tilde{\eta}(1-(2\rho^{2}+\beta)\tilde{\eta}L_{t})\|\nabla L_{t}\|^{2}.$
We complete the proof of the right hand side inequality. The left hand side inequality can be proved similarly. In detail, we can show that:
$\ell(q_{i}(t+1))$ $\geq\ell(q_{i}(t))+\ell^{\prime}(q_{i}(t))(q_{i}(t+1)-q_{i}(t)$
$\geq\ell(q_{i}(t))+\ell^{\prime}(q_{i}(t))\langle y_{i}\nabla f(\mathbf{w}_{t};\mathbf{x}_{i}),\mathbf{w}_{t+1}-\mathbf{w}_{t}\rangle-|\ell^{\prime}(q_{i}(t))|\cdot|\xi[f](\mathbf{w}_{t},\mathbf{w}_{t+1})|.$
Taking the average over all data points, we have
$L_{t+1}\geq L_{t}-\tilde{\eta}(1+\beta\tilde{\eta}L_{t})\|\nabla L_{t}\|^{2}.$
Now we have completed the proof of Lemma A.3. ∎
A.2 Increase of the Parameter Norm
In this section, we demonstrate that the parameter norm, $\rho_{t}$, increases monotonically during the stable phase. We introduce a crucial quantity, $v_{t}$, defined as the inner product of the gradient and the negative weight vector:
$v_{t}\coloneqq\langle\nabla L(\mathbf{w}_{t}),-\mathbf{w}_{t}\rangle.$
This quantity, $v_{t}$, plays a key role in controlling the increase of the parameter norm. Notably, $v_{t}$ appears as the cross term in the expression $\|\mathbf{w}_{t+1}\|^{2}=\|\mathbf{w}_{t}-\tilde{\eta}\nabla L(\mathbf{w}_{t})\|^{2}$. By managing $v_{t}$, we can effectively characterize the increase in the parameter norm.
Recall that our loss function is $\ell(x):=\log(1+e^{-x})$. Inspired by Lyu and Li [2020], we define the following two auxiliary functions for the logistic loss:
$\psi(x)$ $\coloneqq-\log(\ell(x))=-\log\log(1+e^{-x}),\quad x\in\mathbb{R},$ (6)
$\iota(x)$ $\coloneqq\psi^{-1}(x)=-\log(e^{e^{-x}}-1),\quad x\in\mathbb{R}.$
One important remark is that if we change the loss to the exponential loss, both $\psi$ and $\iota$ will be the identity function. Since the logistic loss and the exponential loss have similar tails, our $\psi(x)$ and $\iota(x)$ are close to the identity function, i.e.,
$\psi(x)\approx\iota(x)\approx x,\quad\text{for $x$ large enough.}$.
Then, we have an exponential-loss-like decomposition of $L_{t}$:
$L_{t}=\frac{1}{n}\sum_{i=1}^{n}\ell(q_{i}(t))=\frac{1}{n}\sum_{i=1}^{n}e^{-\psi(q_{i}(t))}.$ (7)
These two functions $\psi,\iota$ will help us to analyze the lower bound of $v_{t}$. First, we list some properties of $\psi$ and $\iota$ here.
Lemma A.4 (Auxiliary functions of $\ell$).
The following claims hold for $\ell$, $\psi$, and $\iota$.
  • $\ell(x)=e^{-\psi(x)}$.
  • $\ell$ is monotonically decreasing, while $\psi$ and $\iota$ are monotonically increasing.
  • $\psi^{\prime}(\iota(x))=\frac{1}{\iota^{\prime}(x)}$;
  • $\psi^{\prime}(x)x$ is increasing for $x\in(0,+\infty)$.
Proof of Lemma A.4.
The first two properties are straightforward. For the third property, we apply chain rule on $\psi(\iota(x))=x$ to get
$\psi^{\prime}(\iota(x))\iota^{\prime}(x)=1.$
For the fourth property, notice that
$\psi^{\prime}(x)x=\frac{x}{(1+e^{x})\log(1+e^{-x})}.$
The denominator is positive and decreasing since
$\frac{d}{dx}\big{[}(1+e^{x})\log(1+e^{-x})\big{]}=e^{x}\log(1+e^{-x})-1\leq e^{x}e^{-x}-1=0.$
Combining this with the fact that $x$ is positive and increasing, we have completed the proof of Lemma A.4. ∎
Besides, we have the following property of $\iota$. This is the key lemma to handle the homogeneous error. Actually, this lemma is another way to show $\iota(x)$ is close to the identity function.
Lemma A.5 (Property of $\iota$).
For every $x\in\mathbb{R}$, we have
$\frac{\iota(x)}{\iota^{\prime}(x)}\geq x+\log\log 2.$
Proof of Lemma A.5.
Recall that
$\iota(x)=-\log(e^{e^{-x}}-1),\quad\iota^{\prime}(x)=\frac{e^{e^{-x}}e^{-x}}{e^{e^{-x}}-1}.$
Let $y=e^{-x}$. We have
$\frac{\iota(x)}{\iota^{\prime}(x)}=\frac{-\log(e^{e^{-x}}-1)(e^{e^{-x}}-1)}{e^{e^{-x}}e^{-x}}=\frac{-\log(e^{y}-1)(e^{y}-1)}{e^{y}y}.$
Define $s(y)\coloneqq\frac{\iota(x)}{\iota^{\prime}(x)}-x-\log\log 2$. Then,
$s(y)$ $=\frac{-\log(e^{y}-1)(e^{y}-1)}{e^{y}y}+\log(y)-\log\log 2,$
$s^{\prime}(y)$ $=\log(e^{y}-1)\cdot\underbrace{\frac{e^{y}-y-1}{e^{y}y^{2}}}_{>0}.$
Note that the sign of $s^{\prime}$ is determined by $\log(e^{y}-1)$. For $0<e^{y}\leq 2$, $s^{\prime}(y)\leq 0$ and $s(y)$ is decreasing; for $e^{y}\geq 2$, $s(y)$ is increasing. Therefore,
$\min_{y\in(0,\infty)}s(y)=s(\log 2)=0.$
Since $x=-\log y$, we have completed the proof of Lemma A.5. ∎
Another important property of $\iota$ is that it can provide a lower bound for $q_{\min}(t)$.
Lemma A.6 ($\iota$ bound $q_{\min}$).
For every $t\geq 0$, we have
$q_{\min}(t)\geq\iota\big{(}-\log(L_{t})-\log n\big{)}.$
Proof of Lemma A.6.
We use 7 to get
$\frac{1}{n}\ell(q_{\min}(t))\leq L_{t}$ $\Rightarrow\frac{1}{n}e^{-\psi(q_{\min}(t))}\leq L_{t}$
$\Rightarrow\psi(q_{\min}(t))\geq-\log n-\log L_{t}$
$\Rightarrow q_{\min}(t)\geq\iota\big{(}-\log(L_{t})-\log n\big{)}.$ by Lemma A.4
Then, we complete the proof of Lemma A.6. ∎
Now, we are ready to give a lower bound of $v_{t}$. The following lemma is an extension of Corollary E.6 in Lyu and Li [2020], where they dealt with a homogeneous model and the exponential loss; we extend this to a non-homogeneous model. The key ingredient is Lemma A.5.
Lemma A.7 (A lower bound of $v_{t}$).
Suppose 1C holds. Consider $v_{t}:=\langle\nabla L(\mathbf{w}_{t}),-\mathbf{w}_{t}\rangle$. If $L_{t}\leq\frac{1}{2ne^{\kappa}}$, then
$v_{t}\geq-L_{t}\log(2ne^{\kappa}L_{t})\geq 0.$
Proof of Lemma A.7.
By definition, we have
$v_{t}$ $:=\langle\nabla L(\mathbf{w}_{t}),-\mathbf{w}_{t}\rangle$
$=-\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i}))y_{i}\langle\nabla f(\mathbf{w}_{t};\mathbf{x}_{i}),\mathbf{w}_{t}\rangle$
$=-\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i}))y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i})-\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i}))\Big{(}y_{i}\langle\nabla f(\mathbf{w}_{t};\mathbf{x}_{i}),\mathbf{w}_{t}\rangle-y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i})\Big{)}$
$\geq-\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i}))y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i})-\kappa L_{t}$
since $|\ell^{\prime}(x)|\leq\ell(x)$ and $|\langle\nabla f(\mathbf{w}_{t};\mathbf{x}_{i}),\mathbf{w}_{t}\rangle-f(\mathbf{w}_{t};\mathbf{x}_{i})|\leq\kappa$ by 1C
$=\frac{1}{n}\sum_{i=1}^{n}e^{-\psi(q_{i}(t))}\psi^{\prime}(q_{i}(t))q_{i}(t)-\kappa L_{t}.\qquad\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}since $\ell(\cdot)=\exp(-\psi(\cdot))$}}$
Applying Lemma A.6 and Lemma A.4, we have
$q_{i}(t)\geq q_{\min}(t)\geq\iota\big{(}-\log(nL_{t})\big{)}:=-\log(e^{nL_{t}}-1)\geq-\log(e^{\frac{1}{2}}-1)\geq 0.$
Then we can apply Lemma A.4 to get
$\psi^{\prime}(q_{i}(t))q_{i}(t)\geq\psi^{\prime}\Big{(}\iota\big{(}-\log(nL_{t})\big{)}\Big{)}\iota\big{(}-\log(nL_{t})\big{)}=\frac{\iota\big{(}-\log(nL_{t})\big{)}}{\iota^{\prime}\big{(}-\log(nL_{t})\big{)}}.$
Invoking Lemma A.5, we have
$\frac{\iota\big{(}-\log(nL_{t})\big{)}}{\iota^{\prime}\big{(}-\log(nL_{t})\big{)}}\geq-\log(nL_{t})+\log\log 2\geq-\log(nL_{t})+\log\log e^{\frac{1}{2}}=-\log(2nL_{t}).$
Putting the above two inequalities together, we have
$\psi^{\prime}(q_{i}(t))q_{i}(t)\geq-\log(2nL_{t}),\quad\text{for every $i=1,\dots,n$}.$
Plugging this back to the bound of $v_{t}$, we get
$v_{t}$ $\geq-\frac{1}{n}\sum_{i=1}^{n}e^{-\psi(q_{i}(t))}\log(2nL_{t})-\kappa L_{t}$
$=-L_{t}\log(2nL_{t})-\kappa L_{t}$
$=-L_{t}\log(2ne^{\kappa}L_{t})\geq 0.$
This completes the proof of Lemma A.7. ∎
Right now, we get a lower bound for $v_{t}$, which is the cross term in the expression of $\|\mathbf{w}_{t+1}\|^{2}$. The next lemma controls the increase of the parameter norm $\rho_{t}$ using $v_{t}$ and $L_{t}$.
Lemma A.8 (The increase of $\rho_{t}$).
Suppose 1A and 1C hold. If $L_{t}\leq\min\Big{\{}\frac{1}{2ne^{\kappa}},\frac{1}{\tilde{\eta}(4\rho^{2}+2\beta)}\Big{\}}$, then
$0\leq 2\tilde{\eta}v_{t}\leq\rho_{t+1}^{2}-\rho_{t}^{2}\leq 2\tilde{\eta}v_{t}\cdot\bigg{(}1-\frac{1}{2\log(2ne^{\kappa}L_{t})}\bigg{)}.$
Proof of Lemma A.8.
By definition, we have
$\rho_{t+1}^{2}-\rho_{t}^{2}$ $=2\tilde{\eta}\langle\nabla L_{t},-\mathbf{w}_{t}\rangle+\tilde{\eta}^{2}\|\nabla L_{t}\|^{2}$
$=2\tilde{\eta}v_{t}+\tilde{\eta}^{2}\|\nabla L_{t}\|^{2}\geq 2\tilde{\eta}v_{t}\geq 0,$
where the last inequality is by Lemma A.7. Besides,
$\rho_{t+1}^{2}-\rho_{t}^{2}$ $=2\tilde{\eta}v_{t}\bigg{(}1+\frac{\tilde{\eta}\|\nabla L_{t}\|^{2}}{2v_{t}}\bigg{)}$
$\leq 2\tilde{\eta}v_{t}\bigg{(}1+\frac{\tilde{\eta}L_{t}^{2}\rho^{2}}{2v_{t}}\bigg{)}$ by $\ell^{\prime}\leq\ell$, 1A, and Lemma A.7
$\leq 2\tilde{\eta}v_{t}\bigg{(}1+\frac{L_{t}}{2v_{t}}\bigg{)}$ by $L_{t}\leq\frac{1}{\tilde{\eta}(4\rho^{2}+2\beta)}$
$\leq 2\tilde{\eta}v_{t}\bigg{(}1-\frac{1}{2\log(2ne^{\kappa}L_{t})}\bigg{)}.$ by Lemma A.7
This completes the proof of Lemma A.8. ∎
A.3 Convergence of the Margin
In this section, we show that the normalized margin of a general predictor converges in the stable phase. Recall that we define the (normalized) margin as
$\bar{\gamma}(\mathbf{w}):=\frac{\min_{i\in[n]}y_{i}f(\mathbf{w};\mathbf{x}_{i})}{\|\mathbf{w}\|_{2}}.$
However, this normalized margin is not a smooth function of $\mathbf{w}$. Instead, we consider a smoothed margin $\gamma^{a}$ as an easy-to-analyze approximator of the normalized margin [Lyu and Li, 2020]
$\gamma^{a}(\mathbf{w}):=\frac{-\log L(\mathbf{w}_{t})}{\|\mathbf{w}\|_{2}}.$ (8)
We see that $\gamma^{a}$ is a good approximator of $\bar{\gamma}$. We can then use $\gamma^{a}$ to analyze the convergence of the normalized margin since they share the same limit (if it exists). While $\gamma^{a}$ is relatively easy to analyze for gradient flow [Lyu and Li, 2020], analyzing that for GD with a large (but fixed) stepsize is hard. To mitigate this issue, we construct another two margins that work well with large stepsize GD following the ideas of Lyu and Li [2020].
Under Assumption 1, we define an auxiliary margin as
$\gamma^{b}(\mathbf{w})\coloneqq\frac{-\log(2ne^{\kappa}L(\mathbf{w}))}{\|\mathbf{w}\|},$ (9)
and a modified margin as
$\gamma^{c}(\mathbf{w}):=\frac{e^{\Phi(L(\mathbf{w}))}}{\|\mathbf{w}\|},\quad\text{where}\ \Phi(x):=\log(-\log(2ne^{\kappa}x))+\frac{1+(4\rho^{2}+2\beta)\tilde{\eta}}{\log(2ne^{\kappa}x)}.$ (10)
These two margins provide a second-order correction when viewing large stepsize GD as a first-order approximation of gradient flow. In the following discussion, we will show that $\bar{\gamma}(\mathbf{w})\approx\gamma^{a}(\mathbf{w})\approx\gamma^{b}(\mathbf{w})\approx\gamma^{c}(\mathbf{w})$. At last, we will use the convergence of $\gamma^{c}(\mathbf{w}_{t})$ to prove $\bar{\gamma}(\mathbf{w}_{t})$ converges.
The following lemma shows that $\bar{\gamma}(\mathbf{w})\approx\gamma^{a}(\mathbf{w})$.
Lemma A.9 (Smoothed margin).
For the smooth margin $\gamma^{a}(\mathbf{w}_{t})$ defined in 8 and the normalized margin $\bar{\gamma}(\mathbf{w}_{t})$ defined in 3, we have
  • When $L_{t}\leq\frac{1}{2n}$, we have
    $q_{\min}(t)\leq-\log L_{t}\leq\log(2n)+q_{\min}(t),$
    and
    $\bar{\gamma}(\mathbf{w}_{t})\leq\gamma^{a}(\mathbf{w}_{t})\leq\bar{\gamma}(\mathbf{w}_{t})+\frac{\log(2n)}{\rho_{t}}.$
  • Assume 1C holds. If $L_{t}\to 0$, then $|\gamma^{a}(\mathbf{w}_{t})-\bar{\gamma}(\mathbf{w}_{t})|\to 0$.
Proof of Lemma A.9.
To prove the first claim, notice that
$L_{t}\leq\frac{1}{2n}\implies\ell(q_{\min}(t))=\log(1+\exp(-q_{\min}(t)))\leq nL_{t}\leq\frac{1}{2}.$
Therefore we have
$e^{-q_{\min}(t)}\leq e^{\frac{1}{2}}-1\leq 1.$
Using $\frac{x}{2}\leq\log(1+x)\leq x$ for $0\leq x\leq 1$, we get
$\frac{1}{2}e^{-q_{\min}(t)}\leq\ell(q_{\min}(t))=\log(1+e^{-q_{\min}(t)})\leq e^{-q_{\min}(t)}.$
Then we can bound $L_{t}$ by
$\frac{1}{2n}e^{-q_{\min}(t)}\leq\frac{1}{n}\ell(q_{\min}(t))\leq L_{t}\leq\ell(q_{\min}(t))\leq e^{-q_{\min}(t)},$
which is equivalent to
$q_{\min}(t)\leq-\log L_{t}\leq\log(2n)+q_{\min}(t).$
Dividing both sides by $\rho_{t}$ proves the second claim:
$\bar{\gamma}(\mathbf{w}_{t}):=\frac{q_{\min}(t)}{\rho_{t}}\leq\gamma^{a}(\mathbf{w}_{t}):=\frac{-\log L_{t}}{\rho_{t}}\leq\bar{\gamma}(\mathbf{w}_{t})+\frac{\log(2n)}{\rho_{t}}=\frac{\log(2n)+q_{\min}(t)}{\rho_{t}}.$
For the last claim, we only need to show that $\rho_{t}\to\infty$. This is because if $L_{t}\to 0$, we have for any $i\in[n]$, $y_{i}f(\mathbf{w}_{t};x_{i})\to\infty$. Using $y_{i}f(\mathbf{w}_{t};x_{i})\leq C_{r,\kappa}\|\mathbf{w}_{t}\|+C_{r}$ from Lemma F.2, we have $\rho_{t}=\|\mathbf{w}_{t}\|_{2}\to\infty$.
Now we have completed the proof of Lemma A.9. ∎
The following lemma shows that $\gamma^{c}(\mathbf{w})\approx\bar{\gamma}(\mathbf{w})$.
Lemma A.10 (Modified and auxiliary margins).
Suppose that Assumption 1 holds. Let $\gamma^{c}(\mathbf{w}_{t})$ be the modified margin as defined in 10 and $\gamma^{b}(\mathbf{w}_{t})$ be the auxiliary margin as defined in 9. If $L_{t}\leq\frac{1}{2ne^{\kappa+2}}$, there exists a constant $c$ such that
$\gamma^{c}(\mathbf{w}_{t})\leq\gamma^{b}(\mathbf{w}_{t})\leq\bar{\gamma}(\mathbf{w}_{t})\leq\bigg{(}1+\frac{c}{\log(1/L(\mathbf{w}_{t}))}\bigg{)}\gamma^{c}(\mathbf{w}_{t}).$
Proof of Lemma A.10.
Step 1. Proof of the first inequalities. Notice that
[…]
$\leq-\log(2ne^{\kappa}L_{t})\quad\text{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}since $L_{t}\leq\frac{1}{2ne^{\kappa}}$, $\exp\bigg{(}\frac{1+(4\rho^{2}+2\beta)\tilde{\eta}}{\log(2ne^{\kappa}L_{t})}\bigg{)}\leq 1$, and $\log(2ne^{\kappa}L_{t})>0$ }}$
$\leq-\log(L_{t})-\log(2n)$
[…]
By the definition of $\gamma^{b}$ and $\gamma^{c}$ as in 10 and 9, we have
$\gamma^{c}(\mathbf{w}_{t}):=\frac{e^{\Phi(L(\mathbf{w}_{t}))}}{\|\mathbf{w}_{t}\|}\leq\frac{-\log(2ne^{\kappa}L_{t})}{\|\mathbf{w}_{t}\|}\eqqcolon\gamma^{b}(\mathbf{w}_{t})\leq\frac{q_{\min}(t)}{\|\mathbf{w}_{t}\|}=:\bar{\gamma}(\mathbf{w}_{t}).$
This completes the proof of the first two inequalities.
Step 2. Proof of the third inequality. First, we have
$\frac{\bar{\gamma}(\mathbf{w}_{t})}{\gamma^{c}(\mathbf{w}_{t})}$ $=\frac{\bar{\gamma}(\mathbf{w}_{t})}{\gamma^{b}(\mathbf{w}_{t})}\cdot\frac{\gamma^{b}(\mathbf{w}_{t})}{\gamma^{c}(\mathbf{w}_{t})}$
$=\frac{q_{\min}(t)}{-\log(2ne^{\kappa}L_{t})}\cdot\exp\bigg{(}\frac{1+(4\rho^{2}+2\beta)\tilde{\eta}}{-\log(2ne^{\kappa}L_{t})}\bigg{)}$ By the definitions of $\bar{\gamma},\gamma^{b},\gamma^{c}$
$\leq\frac{-\log(L_{t})}{-\log(2ne^{\kappa}L_{t})}\cdot\exp\bigg{(}\frac{1+(4\rho^{2}+2\beta)\tilde{\eta}}{-\log(2ne^{\kappa}L_{t})}\bigg{)}$ Since $q_{\min}(t)\leq\log(-L_{t})$ by Lemma A.9
$=\bigg{(}1+\frac{\log(2ne^{\kappa})}{-\log(2ne^{\kappa}L_{t})}\bigg{)}\cdot\exp\bigg{(}\frac{1+(4\rho^{2}+2\beta)\tilde{\eta}}{-\log(2ne^{\kappa}L_{t})}\bigg{)}.$
To simplify the notation, we let $c_{1}\coloneqq 1+(4\rho^{2}+2\beta)\tilde{\eta}$ and $c_{2}=\log(2ne^{\kappa})$. Since $L_{t}\leq\frac{1}{2ne^{\kappa+2}}\Rightarrow-\log(2ne^{\kappa}L_{t})\geq 2>1$, we have
$\frac{1+(4\rho^{2}+2\beta)\tilde{\eta}}{-\log(2ne^{\kappa}L_{t})}=\frac{c_{1}}{-\log(2ne^{\kappa}L_{t})}\leq c_{1}.$
Besides, given $x<c$, we have $e^{x}\leq 1+e^{c}x$. Therefore,
$\exp\bigg{(}\frac{1+(4\rho^{2}+2\beta)\tilde{\eta}}{-\log(2ne^{\kappa}L_{t})}\bigg{)}=\exp\bigg{(}\frac{c_{1}}{-\log(2ne^{\kappa}L_{t})}\bigg{)}\leq 1+\frac{c_{1}\exp(c_{1})}{-\log(2ne^{\kappa}L_{t})}.$
Plugging this into the bound for $\bar{\gamma}(\mathbf{w}_{t})/\gamma^{c}(\mathbf{w}_{t})$, we get
$\frac{\bar{\gamma}(\mathbf{w}_{t})}{\gamma^{c}(\mathbf{w}_{t})}$ $=\bigg{(}1+\frac{c_{2}}{-\log(2ne^{\kappa}L_{t})}\bigg{)}\cdot\exp\bigg{(}\frac{c_{1}}{-\log(2ne^{\kappa}L_{t})}\bigg{)}$
$\leq\bigg{(}1+\frac{c_{2}}{-\log(2ne^{\kappa}L_{t})}\bigg{)}\cdot\bigg{(}1+\frac{\exp(c_{1})c_{1}}{-\log(2ne^{\kappa}L_{t})}\bigg{)}$
$\leq 1+\frac{c_{2}+\exp(c_{1})c_{1}+c_{2}c_{1}\exp(c_{1})}{-\log(2ne^{\kappa}L_{t})}$ Since $-\log(2ne^{\kappa}L_{t})\geq 1$
$=1+\frac{c_{2}+\exp(c_{1})c_{1}+c_{2}c_{1}\exp(c_{1})}{-\log L_{t}-c_{2}}.$
Note that $-\log L_{t}-c_{2}\geq 2>1$. Because $\frac{x}{x-c_{2}}$ is decreasing when $x\geq c_{2}+1$, we have
$\frac{-\log L_{t}}{-\log L_{t}-c_{2}}\leq c_{2}+1\implies\frac{1}{-\log L_{t}-c_{2}}\leq\frac{c_{2}+1}{-\log L_{t}}.$
Plug this inequality into the previous bound for $\bar{\gamma}(\mathbf{w}_{t})/\gamma^{c}(\mathbf{w}_{t})$, we get
$\frac{\bar{\gamma}(\mathbf{w}_{t})}{\gamma^{c}(\mathbf{w}_{t})}\leq 1+\frac{(c_{2}+\exp(c_{1})c_{1}+c_{2}c_{1}\exp(c_{1}))(c_{2}+1)}{-\log L_{t}}.$
Let $c\coloneqq(c_{2}+\exp(c_{1})c_{1}+c_{2}c_{1}\exp(c_{1}))(c_{2}+1)$. This completes the proof of the third inequality, and thus completes the proof of Lemma A.10. ∎
The next lemma shows the convexity of $\Phi$ defined in 10. The convexity will help us analyze the change of $\gamma^{c}(\mathbf{w}_{t})$ in the gradient descent dynamics. Specifically, we are going to use the property that
$\Phi(x)-\Phi(y)\geq\Phi^{\prime}(y)(x-y),\quad\text{for all $x,y$}.$
Lemma A.11 (Convexity of $\Phi$).
The function $\Phi(x)$ defined in 10 is convex for $0<x<\frac{1}{2ne^{2+\kappa}}$.
Proof of Lemma A.11.
Check that
$\Phi^{\prime}(x)=\frac{1-\frac{1}{\log(2ne^{\kappa}x)}(1+(4\rho^{2}+2\beta)\tilde{\eta})}{x\log(2ne^{\kappa}x)},$
and that
$\Phi^{\prime\prime}(x)=\frac{(1+(4\rho^{2}+2\beta)\tilde{\eta})\cdot(2+\log(2ne^{\kappa}x))-\log^{2}(2ne^{\kappa}x)-\log(2ne^{\kappa}x)}{x^{2}\log^{3}(2ne^{\kappa}x)}.$
Note that when $x\leq\frac{1}{2ne^{2+\kappa}}$, we have $\log(2ne^{\kappa}x)\leq-2$, which implies
$02+\log(2ne^{\kappa}x)\leq 0,\ \log(2ne^{\kappa}x)<0,\ \text{and}\ -\log^{2}(2ne^{\kappa}x)-\log(2ne^{\kappa}x)<$ 0 .
Plugging these into the previous equality, we have $\Phi^{\prime\prime}(x)\geq 0$ when $0<x<\frac{1}{2ne^{2+\kappa}}$. Now we have completed the proof of Lemma A.11. ∎
Before we dive into the proof of the monotonic increasing $\gamma^{c}(\mathbf{w}_{t})$, we show that $\gamma^{c}$ is bounded first. The convergence of $\gamma^{c}$ is a direct consequence of the monotonic increasing and the boundedness of $\gamma^{c}$.
Lemma A.12 (An upper bound on $\gamma^{c}$, $\gamma^{b}$, $\gamma^{a}$ and $\gamma^{c}$).
When $L_{t}\leq\Big{\{}\frac{1}{2ne^{\kappa}},\frac{1}{\tilde{\eta}(4\rho^{2}+2\beta)}\Big{\}}$ for $t\geq s$, there exists $B_{0}$ such that
$\gamma^{c}(\mathbf{w}_{t})\leq\gamma^{b}(\mathbf{w}_{t})\leq\gamma^{a}(\mathbf{w}_{t})\leq\bar{\gamma}(\mathbf{w}_{t})+\frac{\log 2n}{\rho_{s}}\leq B_{0}.$
Proof of Lemma A.12.
Apply lemma A.8, we have $\|\mathbf{w}_{t}\|\geq\rho_{t}\geq\rho_{s}$. Then we can apply lemma F.2 and there exists a constant $C_{\rho_{s},\kappa}$ such that for all $i$,
$|y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i})|\leq C_{\rho_{s},\kappa}\|\mathbf{w}_{t}\|.$
Hence,
$\bar{\gamma}(\mathbf{w}_{t})=\frac{\arg\min_{i\in[n]}y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i})}{\|\mathbf{w}_{t}\|}\leq C_{\rho_{s},\kappa}.$
Besides, by Lemma A.9, we have
$\gamma^{a}(\mathbf{w}_{t})\leq\bar{\gamma}(\mathbf{w}_{t})+\frac{\log 2n}{\rho_{t}}\leq C_{\rho_{s},\kappa}+\frac{\log 2n}{\rho_{s}}.$
By Lemma A.10, we have
$\gamma^{c}(\mathbf{w}_{t})\leq\gamma^{b}(\mathbf{w}_{t})\leq\gamma^{a}(\mathbf{w}_{t})\leq C_{\rho_{s},\kappa}+\frac{\log 2n}{\rho_{s}}.$
Let $B_{0}=C_{\rho_{s},\kappa}+\frac{\log 2n}{\rho_{s}}$. Then, we complete the proof of Lemma A.12. ∎
The following lemma is a variant of Proposition 5, item 1, in [Wu et al., 2024] and Lemma 4.8, in [Lyu and Li, 2020]. Before the lemma, we need some auxiliary definitions. let us define
$\bm{\theta}_{t}\coloneqq\frac{\mathbf{w}_{t}}{\|\mathbf{w}_{t}\|},\quad\bm{\nu}_{t}\coloneqq\bm{\theta}_{t}\bm{\theta}_{t}^{\top}(-\nabla L_{t}),\quad\bm{\mu}_{t}\coloneqq\big{(}\mathbf{I}-\bm{\theta}_{t}\bm{\theta}_{t}^{\top}\big{)}(-\nabla L_{t}).$
Therefore, we have
$\|\nabla L_{t}\|^{2}=\|\bm{\nu}_{t}\|^{2}+\|\bm{\mu}_{t}\|^{2}.$
The key point of this decomposition is that we consider the gradient of the loss function as a sum of two orthogonal components. The first component $\bm{\nu}_{t}$ is the component in the direction of $\mathbf{w}_{t}$, and the second component $\bm{\mu}_{t}$ is the component orthogonal to $\mathbf{w}_{t}$. We will show that the modified margin $\gamma^{c}(\mathbf{w}_{t})$ is monotonically increasing. And the increase of $\gamma^{c}(\mathbf{w}_{t})$ is lower bounded by a term that depends on $\|\bm{\mu}_{t}\|^{2}$.
Lemma A.13 (Modified margin is monotonically increasing).
Suppose Assumption 1 holds. If there exists $s$ such that
$L_{s}\leq\min\bigg{\{}\frac{1}{e^{\kappa+2}2n},\ \frac{1}{\tilde{\eta}(4\rho^{2}+2\beta)}\bigg{\}},$
then for $t\geq s$, we have
  • $L_{t+1}\leq L_{t}$.
  • $v_{t}\geq-L_{t}\log(2ne^{\kappa}L_{t})\geq 0$.
  • $\rho_{t+1}^{2}-\rho_{t}^{2}\geq 2\tilde{\eta}v_{t}$.
  • $\log\gamma^{c}(\mathbf{w}_{t+1})-\log\gamma^{c}(\mathbf{w}_{t})\geq\frac{\rho_{t}^{2}}{v_{t}^{2}}\|\bm{\mu}_{t}\|^{2}\log\frac{\rho_{t+1}}{\rho_{t}}$.
As a consequence, $\gamma^{c}(\mathbf{w}_{t})$ admits a finite limit.
Proof of Lemma A.13.
The first claim is by Lemma A.3 and induction. The second and the third claims are consequences of Lemmas A.7 and A.8, respectively. We now prove the last claim. The proof is divided into two steps.
  • In step 1, we constuct an auxilary function $\Psi(x)$ which is close to $-\Phi^{\prime}(x)$ and show that:
    $\Psi(L_{t})(L_{t+1}-L_{t})\leq-\frac{1}{\rho_{t}^{2}}(\rho_{t}^{2}-\rho_{t+1}^{2})\big{(}\frac{1}{2}+\frac{\rho_{t}^{2}}{2v_{t}^{2}}\|\bm{\mu}_{t}\|^{2}\big{)}.$
  • In Step 2, we show that $\Psi(x)\leq-\Phi^{\prime}(x)$ and use this property to get the monotonicity of the modified margin $\gamma^{c}$.
Step 1. Construct $\Psi(x)$.
By Lemma A.3, we have
$L_{t+1}-L_{t}:=L(\mathbf{w}_{t+1})-L(\mathbf{w}_{t})\leq-\tilde{\eta}(1-(2\rho^{2}+\beta)\tilde{\eta}L_{t})\|\nabla L_{t}\|^{2}.$
Multiplying both sides by $02v_{t}\frac{1-\frac{1}{2\log(2ne^{\kappa}L_{t})}}{1-(2\rho^{2}+\beta)\tilde{\eta}L_{t}}>$ 0, we get
$\frac{1-\frac{1}{2\log(2ne^{\kappa}L_{t})}}{1-(2\rho^{2}+\beta)\tilde{\eta}L_{t}}2v_{t}(L_{t+1}-L_{t})\leq-2\tilde{\eta}v_{t}\bigg{(}1-\frac{1}{2\log(2ne^{\kappa}L_{t})}\bigg{)}\|\nabla L_{t}\|^{2}.$
From Lemma A.8 we have
$0\leq\rho_{t+1}^{2}-\rho_{t}^{2}\leq 2\tilde{\eta}v_{t}\Big{(}1-\frac{1}{2\log(2ne^{\kappa}L_{t})}\Big{)}.$
Using the above we get
$\frac{1-\frac{1}{2\log(2ne^{\kappa}L_{t})}}{1-(2\rho^{2}+\beta)\tilde{\eta}L_{t}}2v_{t}(L_{t+1}-L_{t})\leq-(\rho_{t+1}^{2}-\rho_{t}^{2})\|\nabla L_{t}\|^{2}.$ (11)
Recall that
$\|\nabla L_{t}\|^{2}=\|\bm{\nu}_{t}\|^{2}+\|\bm{\mu}_{t}\|^{2}.$
For $\bm{\nu}_{t}$, we have
$\|\bm{\nu}_{t}\|=\frac{1}{\rho_{t}}\langle\mathbf{w}_{t},-\nabla L_{t}\rangle=\frac{v_{t}}{\rho_{t}}.$
Then we can decompose $\|\nabla L_{t}\|^{2}$ as
$\|\nabla L_{t}\|^{2}=\|\bm{\nu}_{t}\|^{2}+\|\bm{\mu}_{t}\|^{2}=\frac{v_{t}^{2}}{\rho_{t}^{2}}+\|\bm{\mu}_{t}\|^{2}.$ (12)
Plugging this into 11 and dividing both two sides by $2v_{t}^{2}$, we have
$\frac{1-\frac{1}{2\log(2ne^{\kappa}L_{t})}}{(1-(2\rho^{2}+\beta)\tilde{\eta}L_{t})v_{t}}(L_{t+1}-L_{t})\leq-\frac{1}{\rho_{t}^{2}}(\rho_{t+1}^{2}-\rho_{t}^{2})\bigg{(}\frac{1}{2}+\frac{\rho_{t}^{2}}{2v_{t}^{2}}\|\bm{\mu}_{t}\|^{2}\bigg{)}.$
By Lemma A.7, we have $v_{t}\geq-L_{t}\log(2ne^{\kappa}L_{t})$. Define
$\Psi(x):=-\frac{1-\frac{1}{2\log(2ne^{\kappa}x)}}{(1-(2\rho^{2}+\beta)\tilde{\eta}x)x\log(2ne^{\kappa}x)}.$
Then we have
$\Psi(L_{t})(L_{t+1}-L_{t})$ $\coloneqq-\frac{1-\frac{1}{2\log(2ne^{\kappa}L_{t})}}{(1-(2\rho^{2}+\beta)\tilde{\eta}L_{t})L_{t}\log(2ne^{\kappa}L_{t})}(L_{t+1}-L_{t})$ (13)
$\leq\frac{1-\frac{1}{2\log(2ne^{\kappa}L_{t})}}{(1-(2\rho^{2}+\beta)\tilde{\eta}L_{t})v_{t}}(L_{t+1}-L_{t})$
$\leq-\frac{1}{\rho_{t}^{2}}(\rho_{t}^{2}-\rho_{t+1}^{2})\bigg{(}\frac{1}{2}+\frac{\rho_{t}^{2}}{2v_{t}^{2}}\|\bm{\mu}_{t}\|^{2}\bigg{)}.$
Step 2. Show $\Psi(x)\leq-\Phi^{\prime}(x)$.
We are going to show that $\Psi(x)\leq-\Phi^{\prime}(x)$. Note that when $0<x\leq\min\Big{\{}\frac{1}{e^{\kappa+2}2n},\frac{1}{\tilde{\eta}(4\rho^{2}+2\beta)}\Big{\}}$, we have $\log(2ne^{\kappa}x)<0$ and $1-(2\rho^{2}+\beta)\tilde{\eta}x)\geq\frac{1}{2}>0$. Therefore, we have
$\Psi(x)=\underbrace{\frac{1-\frac{1}{2\log(2ne^{\kappa}x)}}{1-(2\rho^{2}+\beta)\tilde{\eta}x}}_{\eqqcolon J>0}\cdot\underbrace{\frac{-1}{x\log(2ne^{\kappa}x)}}_{>0}.$ (14)
To get an upper bound of $\Psi(x)$, we just need an upper bound of $J$. Let $a:=\frac{-1}{2\log(2ne^{\kappa}x)}>0$ and $b:=(2\rho^{2}+\beta)\tilde{\eta}x\in(0,1/2]$. Then we invoke Lemma F.4 to get
$J:=\frac{1+a}{1-b}\leq 1+2a+2b=1-\frac{1}{\log(2ne^{\kappa}x)}+(4\rho^{2}+2\beta)\tilde{\eta}x.$
Recall that $x\leq\frac{1}{e^{\kappa+2}2n}\leq\frac{1}{2ne^{\kappa}}$ and $2ne^{\kappa}\geq 1$. Then we apply Lemma F.3 to get $x\leq\frac{-1}{\log(2ne^{\kappa}x)}$. Plugging this into the bound of $J$, we get
$J\leq 1-\frac{1}{\log(2ne^{\kappa}x)}+(4\rho^{2}+2\beta)\tilde{\eta}x\leq 1-\frac{1}{\log(2ne^{\kappa}x)}(1+(4\rho^{2}+2\beta)\tilde{\eta}).$ (15)
Plugging (15) into (14), we have
$\Psi(x)=J\cdot\frac{-1}{x\log(2ne^{\kappa}x)}\leq-\frac{1-\frac{1}{\log(2ne^{\kappa}x)}(1+(4\rho^{2}+2\beta)\tilde{\eta})}{x\log(2ne^{\kappa}x)}=-\Phi^{\prime}(x),$
which verifies that $\Psi(x)\leq-\Phi^{\prime}(x)$. By this and 13, we have
$\Phi^{\prime}(L_{t})(L_{t+1}-L_{t})+\varphi^{\prime}(\rho_{t}^{2})(\rho_{t+1}^{2}-\rho_{t}^{2})\bigg{(}\frac{1}{2}+\frac{\rho_{t}^{2}}{2v_{t}^{2}}\|\bm{\mu}_{t}\|^{2}\bigg{)}\geq 0,$
where $\varphi(x)=-\log x=\log(1/x)$. Recall that for $0<x\leq\frac{1}{2ne^{\kappa+2}}$, $\Phi(x)$ is convex by Lemma A.11. By convexity of $\varphi$ and $\Phi$, we have
$\Phi(L_{t+1})-\Phi(L_{t})+\bigg{(}\log\frac{1}{\rho_{t+1}^{2}}-\log\frac{1}{\rho_{t}^{2}}\bigg{)}\bigg{(}\frac{1}{2}+\frac{\rho_{t}^{2}}{2v_{t}^{2}}\|\bm{\mu}_{t}\|^{2}\bigg{)}\geq 0.$
By the definition of $\gamma^{c}$ in 10, this can be rewritten as
$\log\gamma^{c}(\mathbf{w}_{t+1})-\log\gamma^{c}(\mathbf{w}_{t})$ $=(\Phi(L_{t+1})-\Phi(L_{t}))+\bigg{(}\log\frac{1}{\rho_{t+1}}-\log\frac{1}{\rho_{t}}\bigg{)}$
$\geq-\bigg{(}\log\frac{1}{\rho_{t+1}^{2}}-\log\frac{1}{\rho_{t}^{2}}\bigg{)}\frac{\rho_{t}^{2}}{2v_{t}^{2}}\|\bm{\mu}_{t}\|^{2}$
$=\frac{\rho_{t}^{2}}{v_{t}^{2}}\|\bm{\mu}_{t}\|^{2}\log\frac{\rho_{t+1}}{\rho_{t}}$
$\geq 0,$
where the last inequality is because of Lemma A.8. We have shown that $\gamma^{c}(\mathbf{w}_{t})$ is monotonically increasing. By Lemma A.12, $\gamma^{c}(\mathbf{w}_{t})$ is bounded. Therefore $\gamma^{c}(\mathbf{w}_{t})$ admits a finite limit. This completes the proof of Lemma A.13. ∎
A.4 Sharp rates of Loss and Parameter Norm
Right now, we have already proved that $\gamma^{c}(\mathbf{w}_{t})$ is monotonically increasing and bounded, which indicates $\gamma^{c}(\mathbf{w}_{t})$ converges. However, if we want to show that $\bar{\gamma}(\mathbf{w}_{t})$ converges, we still need to verify that $L_{t}\to 0$, which is the crucial condition for $\gamma^{c}(\mathbf{w}_{t}),\gamma^{b}(\mathbf{w}_{t}),\gamma^{a}(\mathbf{w}_{t})$, and $\bar{\gamma}(\mathbf{w}_{t})$ to share the same limit, by Lemma A.9 and Lemma A.10.
Fortunately, with the monotonicity of $\gamma^{c}(\mathbf{w}_{t})$, we can prove that $L_{t}$ converges to zero and even characterize the rate of $L_{t}$.
Lemma A.14 (Rate of $L_{t}$ in general model).
Suppose Assumption 1 holds. If there is an $s$ such that
$L(\mathbf{w}_{s})\leq\min\big{\{}\frac{1}{e^{\kappa+2}2n},\frac{1}{\tilde{\eta}(4\rho^{2}+2\beta)}\big{\}},$
then for every $t\geq s$ we have
$\frac{1}{\frac{1}{L(\mathbf{w}_{s})}+3\tilde{\eta}\rho^{2}(t-s)}\leq L(\mathbf{w}_{t})\leq\frac{2}{(t-s)\tilde{\eta}\gamma^{c}(\mathbf{w}_{s})^{2}}.$
That is, $L(\mathbf{w}_{t})=\Theta(1/t)\to 0$ as $t\to\infty$.
Proof of Lemma A.14.
By Lemma A.3 and 12 in the proof of Lemma A.13, we know $L_{t}$ is decreasing and
$L_{t+1}-L_{t}\leq-\frac{\tilde{\eta}}{2}\|\nabla L_{t}\|^{2}\leq-\frac{\tilde{\eta}}{2}\|\bm{\nu}_{t}\|_{2}^{2}\leq-\frac{\tilde{\eta}}{2}\frac{v_{t}^{2}}{\rho_{t}^{2}}.$ (16)
We will establish an upper bound for $\rho_{t}$ first. Note that $\gamma^{c}(\mathbf{w}_{t})$ is increasing for $t\geq s$ by Lemma A.13 and $\gamma^{b}(\mathbf{w}_{t})\geq\gamma^{c}(\mathbf{w}_{t})$ by Lemma A.10. By 9, we have
$\rho_{t}=\frac{-\log(2ne^{\kappa}L_{t})}{\gamma^{b}(\mathbf{w}_{t})}\leq\frac{-\log(2ne^{\kappa}L_{t})}{\gamma^{c}(\mathbf{w}_{t})}\leq\frac{-\log(2ne^{\kappa}L_{t})}{\gamma^{c}(\mathbf{w}_{s})}.$
Combining this with Lemma A.7, we have
$\frac{v_{t}}{\rho_{t}}\geq\frac{-L_{t}\log(2ne^{\kappa}L_{t})}{\frac{-\log(2ne^{\kappa}L_{t})}{\gamma^{c}(\mathbf{w}_{s})}}=L_{t}\gamma^{c}(\mathbf{w}_{s}).$
Plugging this into (16), we have
$L_{t+1}-L_{t}\leq-\frac{\tilde{\eta}}{2}L_{t}^{2}\gamma^{c}(\mathbf{w}_{s})^{2},$
which implies
$\frac{\tilde{\eta}\gamma(\mathbf{w}_{s})^{2}}{2}$ $\leq\frac{L_{t}-L_{t+1}}{L_{t}^{2}}$
$\leq\frac{L_{t}-L_{t+1}}{L_{t}L_{t+1}}$ Since $L_{t+1}\leq L_{t}$
$=\frac{1}{L_{t+1}}-\frac{1}{L_{t}},\quad t\geq s.$
Telescoping the sum from $s$ to $t$, we have
$(t-s)\frac{\tilde{\eta}\gamma^{c}(\mathbf{w}_{s})^{2}}{2}\leq\frac{1}{L_{t}}-\frac{1}{L_{s}}\leq\frac{1}{L_{t}}.$
Therefore we have
$L_{t}\leq\frac{2}{(t-s)\tilde{\eta}\gamma^{c}(\mathbf{w}_{s})^{2}}.$
Next we show the lower bound on the risk. By Lemma A.3 we have
$L_{t+1}-L_{t}\geq-\tilde{\eta}(1+\beta\tilde{\eta}L_{t})\|\nabla L_{t}\|^{2}\geq\tilde{-}\frac{3}{2}\eta\|\nabla L_{t}\|^{2}.$
Observe that under 1A,
$\|\nabla L_{t}\|=\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(q_{i}(t))y_{i}\nabla f(\mathbf{w}_{t};\mathbf{x}_{i})\bigg{\|}\leq\rho L_{t}.$
Then we have
$L_{t+1}-L_{t}\geq-\tilde{\eta}\frac{3}{2}\rho^{2}L_{t}^{2},\quad t\geq s.$
Let $\tilde{L}_{t}:=\frac{3\tilde{\eta}\rho^{2}}{2}L_{t}$, we have $\tilde{L}_{s}\leq\frac{3\tilde{\eta}\rho^{2}}{2}\frac{1}{\tilde{\eta}(4\rho^{2}+2\beta)}\leq\frac{3}{8}\leq\frac{1}{2}$. Furthermore, since $L_{t}$ decreases monotonically, $\tilde{L}_{t}\leq\tilde{L}_{s}\leq\frac{1}{2}$. The inequality becomes
$\tilde{L}_{t+1}-\tilde{L}_{t}\geq-\tilde{L}_{t}^{2}.$
Therefore, let $c=\frac{1}{\tilde{L}_{s}}$ and apply Lemma F.1, we have for any $t\geq s$,
$\tilde{L}_{t}\geq\frac{1}{c+2(t-s)}.$
This is equivalent to
$L_{t}\geq\frac{1}{\frac{1}{L_{s}}+3\tilde{\eta}\rho^{2}(t-s)}.$
We have completed the proof of Lemma A.14. ∎
Furthermore, we can characterize the order of $\rho_{t}$ in the stable phase.
Lemma A.15 (Order of $\rho_{t}$ in general model).
Suppose Assumption 1 holds. If there is $s$ such that
$L(\mathbf{w}_{s})\leq\min\bigg{\{}\frac{1}{e^{\kappa+2}2n},\ \frac{1}{\tilde{\eta}(4\rho^{2}+2\beta)}\bigg{\}},$
then for $t\geq s$ we have
$\rho_{t}=\Theta(\log(t)).$
Proof of Lemma A.15.
Note that $\gamma^{c}(\mathbf{w}_{t})$ is increasing for $t\geq s$ by Lemma A.13 and $\gamma^{b}(\mathbf{w}_{t})\geq\gamma^{c}(\mathbf{w}_{t})$ by Lemma A.10. Therefore,
$\rho_{t}\leq\frac{-\log(2ne^{\kappa}L_{t})}{\gamma^{b}(\mathbf{w}_{t})}\leq\frac{-\log(2ne^{\kappa}L_{t})}{\gamma^{c}(\mathbf{w}_{t})}\leq\frac{-\log(2ne^{\kappa}L_{t})}{\gamma^{c}(\mathbf{w}_{s})}.$
Combining this with Lemma A.14, we have
$\rho_{t}\leq\frac{\log\frac{1/L(\mathbf{w}_{s})+3\tilde{\eta}\rho^{2}(t-s)}{2ne^{\kappa}}}{\gamma^{c}(\mathbf{w}_{s})}=\mathcal{O}(\log(\tilde{\eta}t)).$
Besides, we have $q_{\min}\geq\iota(\log\frac{1}{L}-\log n)$ by Lemma A.9 and $q_{\min}\leq B_{0}\rho_{t}$ by Lemma A.12. Therefore we have
$\rho_{t}\geq\frac{\iota(\log\frac{1}{nL_{t}})}{B_{0}}\geq\frac{\log\frac{1}{nL_{t}}}{2B_{0}}\geq\frac{\log\frac{(t-s)\tilde{\eta}\gamma^{c}(\mathbf{w}_{s})^{2}}{2n}}{2B_{0}}=\Omega(\log(t)),$
where the second inequality is because for $\iota(x)$ defined in (6), $\iota(x)\geq\frac{x}{2}$ for $x\geq 0.6$, and the third inequality is by Lemma A.14. Combining them, we get
$\rho_{t}=\Theta(\log(t)).$
This completes the proof of Lemma A.15. ∎
A.5 Proof of Theorem 2.2
Proof of Theorem 2.2.
We prove the items one by one.
  • The monotonicity of $L_{t}$ comes from the result of Lemma A.3 directly.
  • Item 1 is due to Lemma A.14 .
  • For item 2, the monotonicity of $\rho_{t}$ comes from the result of Lemma A.8 and the order is due to Lemma A.15.
[…]
Appendix B EoS Phase Analysis
In this section, we focus on the linearly separable case, that is, we work under Assumption 3. We mainly follow the idea of [Wu et al., 2024] for the proof. In detail, we consider a comparator
$\mathbf{u}:=\mathbf{u}_{1}+\mathbf{u}_{2},$
where where
$\mathbf{u}_{1}:=\begin{pmatrix}\mathbf{u}_{1}^{(1)}\\ \vdots\\ \mathbf{u}_{1}^{(m)}\end{pmatrix},\ \ \text{with}\ \ \mathbf{u}_{1}^{(j)}:=a_{j}\frac{\log(\gamma^{2}\eta T)+\kappa}{\alpha\gamma}\cdot\mathbf{w}_{*},\quad j=1,\ldots,m,$ (17)
and
$\mathbf{u}_{2}:=\begin{pmatrix}\mathbf{u}_{2}^{(1)}\\ \vdots\\ \mathbf{u}_{2}^{(m)}\end{pmatrix},\ \ \text{with}\ \ \mathbf{u}_{2}^{(j)}:=a_{j}\frac{\eta}{2\gamma}\cdot\mathbf{w}_{*},\quad j=1,\ldots,m.$ (18)
Consider the following decomposition,
[…]
We aim to prove $I_{1}(\mathbf{w}_{t})\leq\frac{1}{T}-L(\mathbf{w}_{t})$ and $I_{2}(\mathbf{w}_{t})\leq 0$. Then we can get a bound for the average loss by telescope summing the decomposition. Here we also introduced the following vector $\bar{\mathbf{w}}_{*}$:
$\bar{\mathbf{w}}_{*}:=\begin{pmatrix}a_{1}\mathbf{w}_{*}\\ \vdots\\ a_{n}\mathbf{w}_{*}\end{pmatrix}$
We can observe that $\mathbf{u}_{1}=\frac{\log(\gamma^{2}\eta T)+\kappa}{\alpha\gamma}\bar{\mathbf{w}}_{*}$ and $\mathbf{u}_{2}=\frac{\eta}{2\gamma}\bar{\mathbf{w}}_{*}$.
Lemma B.1 (A bound on $I_{1}(\mathbf{w})$ in the EoS phase).
For $\mathbf{u}_{1}$ defined in (17), we have
$I_{1}(\mathbf{w}):=\langle\nabla L(\mathbf{w}),\mathbf{u}_{1}-\mathbf{w}\rangle\leq\frac{1}{\gamma^{2}\eta T}-L(\mathbf{w}).$
Proof of Lemma B.1.
Since $L$ is averaged over the individual losses incurred at the data $(\mathbf{x}_{i},y_{i})_{i=1}^{n}$ and gradient is a linear operator, it suffices to prove the claim assuming there is only a single data point $(\mathbf{x},y)$. Then by Assumption 3, we have
$\langle y\mathbf{x},\mathbf{w}_{*}\rangle\geq\gamma>0.$
Then the loss becomes
$L(\mathbf{w})=\ell(yf(\mathbf{w};\mathbf{x}))=\ell\biggl{(}y\frac{1}{m}\sum_{j=1}^{m}a_{j}\phi(\mathbf{x}^{\top}\mathbf{w}^{(j)})\biggr{)}.$
Now we expand $I_{1}(\mathbf{w})$:
$I_{1}(\mathbf{w})$ $:=\langle\nabla L(\mathbf{w}),\mathbf{u}_{1}-\mathbf{w}\rangle$
$=\ell^{\prime}\big{(}yf(\mathbf{w};\mathbf{x})\big{)}\langle y\nabla f(\mathbf{w};\mathbf{x}),\mathbf{u}_{1}-\mathbf{w}\rangle$
$=\ell^{\prime}\bigl{(}yf(\mathbf{w};\mathbf{x})\bigr{)}\frac{1}{m}\sum_{k=1}^{m}a_{k}y\phi^{\prime}(x^{\top}\mathbf{w}^{(k)})\mathbf{x}^{\top}(\mathbf{u}_{1}^{(k)}-\mathbf{w}^{(k)})$
$=\ell^{\prime}\bigl{(}yf(\mathbf{w};\mathbf{x})\bigr{)}\Biggl{[}\underbrace{\frac{1}{m}\sum_{k=1}^{m}a_{k}y\Bigl{(}\phi^{\prime}(\mathbf{x}^{\top}\mathbf{w}^{(k)})\mathbf{x}^{\top}\mathbf{u}_{1}^{(k)}+\phi(\mathbf{x}^{\top}\mathbf{w}^{(k)})-\phi^{\prime}(\mathbf{x}^{\top}\mathbf{w}^{(k)})\mathbf{x}^{\top}\mathbf{w}^{(k)}\Bigr{)}}_{=:J_{1}}$
$\hskip 227.62204pt-\underbrace{\frac{1}{m}\sum_{k=1}^{m}a_{k}y\phi(\mathbf{x}^{\top}\mathbf{w}^{(k)})}_{=:J_{2}}\Biggr{]}.$ (19)
By definition we have $J_{2}=yf(\mathbf{w};\mathbf{x}).$ As for $J_{1}$, using $\phi^{\prime}\geq\alpha$ and $a_{k}yx^{\top}\mathbf{u}_{1}^{(k)}\geq 0$ by Assumption 3, we have
$J_{1}$ $:=\frac{1}{m}\sum_{k=1}^{m}a_{k}y\Bigl{(}\phi^{\prime}(\mathbf{x}^{\top}\mathbf{w}^{(k)})\mathbf{x}^{\top}\mathbf{u}_{1}^{(k)}+\phi(\mathbf{x}^{\top}\mathbf{w}^{(k)})-\phi^{\prime}(\mathbf{x}^{\top}\mathbf{w}^{(k)})\mathbf{x}^{\top}\mathbf{w}^{(k)}\Bigr{)}$
$\geq\frac{1}{m}\sum_{k=1}^{m}a_{k}\alpha y\mathbf{x}^{\top}\mathbf{u}_{1}^{(k)}+\frac{1}{m}\sum_{k=1}^{m}a_{k}y\Bigl{(}\phi(\mathbf{x}^{\top}\mathbf{w}^{(k)})-\phi^{\prime}(\mathbf{x}^{\top}\mathbf{w}^{(k)})\mathbf{x}^{\top}\mathbf{w}^{(k)}\Bigr{)}$
$\geq\frac{1}{m}\sum_{k=1}^{m}a_{k}^{2}\alpha\frac{\log(\gamma^{2}\eta T)+\kappa}{\alpha\gamma}y\mathbf{x}^{\top}\mathbf{w}_{*}-\frac{1}{m}\sum_{k=1}^{m}|a_{k}|\kappa$
since $|\phi(\mathbf{x}^{\top}\mathbf{w}^{(k)})-\phi^{\prime}(\mathbf{x}^{\top}\mathbf{w}^{(k)})\mathbf{x}^{\top}\mathbf{w}^{(k)}|\leq\kappa$ by 2C
$\geq\log(\gamma^{2}\eta T)+\kappa-\kappa$
since $y\mathbf{x}^{\top}\mathbf{w}_{*}\geq\gamma$ and $\sum_{k=1}^{m}a_{k}^{2}=m$
$=\log(\gamma^{2}\eta T).$ (20)
Plugging in $J_{2}=yf(\mathbf{w};\mathbf{x})$ and (B) into (B), we get
$I_{1}(\mathbf{w})$ $=\langle\nabla L(\mathbf{w}),\mathbf{u}_{1}-\mathbf{w}\rangle=\ell^{\prime}\big{(}yf(\mathbf{w};\mathbf{x})\big{)}(J_{1}-J_{2})$
$\leq\ell^{\prime}\bigl{(}yf(\mathbf{w};\mathbf{x})\bigr{)}\Bigl{[}\log(\gamma^{2}\eta T)-yf(\mathbf{w};\mathbf{x})\Bigr{]}$ since $\ell^{\prime}<0$
$\leq\ell(\log(\gamma^{2}\eta T))-\ell(yf(\mathbf{w};\mathbf{x}))$ since $\ell$ is convex
$\leq\frac{1}{\gamma^{2}\eta T}-L(\mathbf{w}).$
where in the last inequality, we use $\ell(x)\leq\exp(-x)$ and we only consider a single data point. This completes the proof of Lemma B.1. ∎
Lemma B.2 (A bound on $I_{2}(\mathbf{w})$ in EoS).
For $\mathbf{u}_{2}$ defined in (18), for every $\mathbf{w}$,
$I_{2}(\mathbf{w}):=2\langle\nabla L(\mathbf{w}),\mathbf{u}_{2}\rangle+m\eta\|\nabla L(\mathbf{w})\|^{2}\leq 0.$
Proof of Lemma B.2.
For simplicity, we define
$g_{i}(\mathbf{w}^{(j)})\coloneqq\ell^{\prime}(y_{i}f(\mathbf{w};\mathbf{x}_{i}))\phi^{\prime}(\mathbf{x}_{i}^{\top}\mathbf{w}^{(j)}).$
Note that $-1\leq\ell^{\prime}(\cdot)\leq 0$ and $0<\alpha\leq\phi^{\prime}(\cdot)\leq 1$, we have
$-1\leq g_{i}(\mathbf{w}^{(j)})\leq 0.$
Under this notation, we have
$\frac{\partial L(\mathbf{w})}{\partial\mathbf{w}_{i}}$ $=\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}\big{(}y_{i}f(\mathbf{w};\mathbf{x}_{i})\big{)}y_{i}a_{j}m^{-1}\phi^{\prime}\big{(}\mathbf{x}_{i}^{\top}\mathbf{w}^{(j)}\big{)}\mathbf{x}_{i}$
$=\frac{1}{n}\sum_{i=1}^{n}g_{i}\big{(}\mathbf{w}^{(j)}\big{)}a_{j}m^{-1}y_{i}\mathbf{x}_{i}.$
So we have
$I_{2}(\mathbf{w})$ $:=2\langle\nabla L(\mathbf{w}),\mathbf{u}_{2}\rangle+m\eta\|\nabla L(\mathbf{w})\|^{2}$
$=\frac{1}{m}\sum_{j=1}^{m}\Biggl{[}\frac{2}{n}\sum_{i=1}^{n}g_{i}(\mathbf{w}^{(j)})a_{j}y_{i}\cdot\mathbf{x}_{i}^{\top}\mathbf{u}_{2}^{(j)}+\eta\biggl{\|}\frac{1}{n}\sum_{i=1}^{n}g_{i}(\mathbf{w}^{(j)})a_{j}y_{i}\mathbf{x}_{i}\biggr{\|}^{2}\Biggr{]}.$
For the term inside the bracket, we have
$\frac{2}{n}\sum_{i=1}^{n}g_{i}(\mathbf{w}^{(j)})a_{j}y_{i}\cdot\mathbf{x}_{i}^{\top}\mathbf{u}_{2}^{(j)}+\eta\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}g_{i}(\mathbf{w}^{(j)})a_{j}y_{i}\mathbf{x}_{i}\bigg{\|}^{2}$
$=$ $\frac{2}{n}\sum_{i=1}^{n}g_{i}(\mathbf{w}^{(j)})a_{j}y_{i}\cdot\mathbf{x}_{i}^{\top}\frac{\eta}{2\gamma}a_{j}\mathbf{w}_{*}+\eta\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}g_{i}(\mathbf{w}^{(j)})a_{j}y_{i}\mathbf{x}_{i}\bigg{\|}^{2}$ since $\mathbf{u}_{2}^{(j)}:=\frac{\eta a_{j}}{2\gamma}\mathbf{w}_{*}$ by (18)
$\leq$ $\frac{2}{n}\sum_{i=1}^{n}g_{i}(\mathbf{w}^{(j)})a_{j}^{2}\frac{\eta}{2\gamma}\gamma+\eta\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}g_{i}(\mathbf{w}^{(j)})a_{j}y_{i}\mathbf{x}_{i}\bigg{\|}^{2}$ since $g_{i}(\cdot)\leq 0$ and $y_{i}x_{i}^{\top}\mathbf{w}_{*}\geq\gamma$
$=$ $\eta\bigg{(}\frac{1}{n}\sum_{i=1}^{n}g_{i}(\mathbf{w}^{(j)})+\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}g_{i}(\mathbf{w}^{(j)})y_{i}\mathbf{x}_{i}\bigg{\|}^{2}\bigg{)}$ since $a_{j}^{2}=1$
$\leq$ $\eta\bigg{(}\frac{1}{n}\sum_{i=1}^{n}g_{i}(\mathbf{w}^{(j)})+\frac{1}{n}\sum_{i=1}^{n}g_{i}^{2}(\mathbf{w}^{(j)})\bigg{)}$ since $|g_{i}(\cdot)|\leq 1$ and $\|y\mathbf{x}\|\leq 1$
$\leq$ $0$ since $-1\leq g_{i}(\cdot)\leq 0$
Hence, we prove that $I_{2}(\mathbf{w})\leq 0$. This completes the proof of Lemma B.2. ∎
Theorem B.3 (A split optimization bound).
For every $\eta>0$ and $\mathbf{u}=\mathbf{u}_{1}+\mathbf{u}_{2}$ such that
$\mathbf{u}_{1}:=\begin{pmatrix}\mathbf{u}_{1}^{(1)}\\ \vdots\\ \mathbf{u}_{1}^{(m)}\end{pmatrix},\ \ \text{with}\ \ \mathbf{u}_{1}^{(j)}:=a_{j}\frac{\log(\gamma^{2}\eta t)+\kappa}{\alpha\gamma}\cdot\mathbf{w}_{*},\quad j=1,\ldots,m,$
and
$\mathbf{u}_{2}:=\begin{pmatrix}\mathbf{u}_{2}^{(1)}\\ \vdots\\ \mathbf{u}_{2}^{(m)}\end{pmatrix},\ \ \text{with}\ \ \mathbf{u}_{2}^{(j)}:=a_{j}\frac{\eta}{2\gamma}\cdot\mathbf{w}_{*},\quad j=1,\ldots,m.$
we have:
$\frac{\|\mathbf{w}_{T}-\mathbf{u}\|^{2}}{2m\eta T}+\frac{1}{T}\sum_{k=0}^{T-1}L\bigl{(}\mathbf{w}^{(k)}\bigr{)}\leq\frac{1+8\log^{2}(\gamma^{2}\eta T)/\alpha^{2}+8\kappa^{2}/\alpha^{2}+\eta^{2}}{\gamma^{2}\eta T}+\frac{\|\mathbf{w}_{0}\|^{2}}{m\eta T},$
for all $T$.
Proof of Theorem B.3.
By Lemma B.1 and Lemma B.2, we have
$\|\mathbf{w}_{t+1}-\mathbf{u}\|^{2}$ $=\|\mathbf{w}_{t}-\mathbf{u}\|^{2}+2m\eta I_{1}(\mathbf{w}_{t})+\eta mI_{2}(\mathbf{w}_{t})$
$\leq\|\mathbf{w}_{t}-\mathbf{u}\|^{2}+2m\eta I_{1}(\mathbf{w}_{t})$
$\leq\|\mathbf{w}_{t}-\mathbf{u}\|^{2}+2m\eta\bigg{(}\frac{1}{\gamma^{2}\eta T}-L(\mathbf{w}_{t})\bigg{)}.$
Telescoping the sum, we get
$\frac{\|\mathbf{w}_{T}-\mathbf{u}\|^{2}}{2m\eta}+\sum_{t=0}^{T-1}L(\mathbf{w}_{t})\leq 1+\frac{\|\mathbf{w}_{0}-\mathbf{u}\|^{2}}{2m\eta}.$
By 17and 18, we have
$\|\mathbf{w}_{0}-\mathbf{u}\|^{2}$ $\leq 2\|\mathbf{w}_{0}\|^{2}+2\|\mathbf{u}\|^{2}$
$\leq 2\|\mathbf{w}_{0}\|_{2}^{2}+4\|\mathbf{u}_{1}\|^{2}+4\|\mathbf{u}_{2}\|^{2}$
$=2\|\mathbf{w}_{0}\|_{2}^{2}+\frac{8m\log(\gamma^{2}\eta T)^{2}+8m\kappa^{2}}{\alpha^{2}\gamma^{2}}+\frac{m\eta^{2}}{\gamma^{2}},$
which implies that
$\frac{\|\mathbf{w}_{T}-\mathbf{u}\|^{2}}{2m\eta t}+\frac{1}{T}\sum_{k=0}^{T-1}L\bigl{(}\mathbf{w}^{(k)}\bigr{)}\leq\frac{1+8\log^{2}(\gamma^{2}\eta T)/\alpha^{2}+8\kappa^{2}/\alpha^{2}+\eta^{2}}{\gamma^{2}\eta T}+\frac{\|\mathbf{w}_{0}\|^{2}}{m\eta T}.$
We complete the proof of Theorem B.3. ∎
B.1 Proof of Theorem 3.2
Proof of Theorem 3.2.
By Theorem B.3, we have
$\frac{1}{T}\sum_{k=0}^{T-1}L\bigl{(}\mathbf{w}^{(k)}\bigr{)}\leq\frac{1+8\log^{2}(\gamma^{2}\eta T)/\alpha^{2}+8\kappa^{2}/\alpha^{2}+\eta^{2}}{\gamma^{2}\eta T}+\frac{\|\mathbf{w}_{0}\|^{2}}{m\eta T}.$
This completes the proof. ∎
Appendix C Phase Transition Analysis
In this section, we will analyze the phase transition. In detail, we follow the idea of [Wu et al., 2024] and apply the perceptron argument [Novikoff, 1962] to locate the phase transition time. Compare to the previous EoS phase analysis, we need an extra assumption on the smoothness of the activation function, which is the 2B.
To proceed, let us define the following quantities for the GD process:
$G(\mathbf{w}):=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{1+\exp\big{(}y_{i}f(\mathbf{w};\mathbf{x}_{i})\big{)}},\quad F(\mathbf{w}):=\frac{1}{n}\sum_{i=1}^{n}\exp\big{(}-y_{i}f(\mathbf{w};\mathbf{x}_{i})\big{)}.$
Due to the self-boundedness of the logistic function, we can show that $G(\mathbf{w}),L(\mathbf{w}),F(\mathbf{w})$ are equivalent in the following sense.
Lemma C.1 (Equivalence of $G,L,F$).
  • 1.
    $G(\mathbf{w})\leq L(\mathbf{w})\leq F(\mathbf{w})$.
  • 2.
    $\alpha\gamma G(\mathbf{w})\leq\sqrt{m}\|\nabla L(\mathbf{w})\|\leq G(\mathbf{w})$.
  • 3.
    If $G(\mathbf{w})\leq\frac{1}{2n}$, then $F(\mathbf{w})\leq 2G(\mathbf{w})$.
Proof of Lemma C.1.
The first claim is by the property of the logistic loss. For the second one,
$\|\nabla L(\mathbf{w})\|^{2}$ $=\sum_{j=1}^{m}\Biggl{\|}\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}f(\mathbf{w};\mathbf{x}_{i}))\cdot y_{i}\cdot a_{j}m^{-1}\phi(\mathbf{x}_{i}^{\top}\mathbf{w}^{(j)})\mathbf{x}_{i}\Biggr{\|}_{2}^{2}$
$\leq\sum_{j=1}^{m}\Biggl{(}\frac{1}{n}\sum_{i=1}^{n}\ell^{\prime}(y_{i}f(\mathbf{w};\mathbf{x}_{i}))\cdot m^{-1}\Biggr{)}^{2}$ since $\|y_{i}a_{j}\phi(\mathbf{x}_{i}^{\top}\mathbf{w}^{(j)})\mathbf{x}_{i}\|\leq 1$
$=\frac{1}{m}G^{2}(\mathbf{w}).$
Besides, we have
$\sqrt{m}\|\nabla L(\mathbf{w})\|$ $\geq\langle-\nabla L(\mathbf{w}),\bar{\mathbf{w}}_{*}\rangle$ since $\|\bar{\mathbf{w}}_{*}\|\leq\sqrt{m}$
$=-\frac{1}{nm}\sum_{i=1}^{n}\sum_{j=1}^{m}\ell^{\prime}(y_{i}f(\mathbf{w};\mathbf{x}_{i}))y_{i}\phi^{\prime}(\mathbf{x}_{i}^{\top}\mathbf{w}_{*})\mathbf{x}_{i}^{\top}\mathbf{w}_{*}$
$\geq\alpha\gamma\frac{1}{n}\sum_{i=1}^{n}\frac{1}{1+\exp\big{(}y_{i}f(\mathbf{w};\mathbf{x}_{i})\big{)}}$ since $\phi^{\prime}\geq\alpha$ and $y_{i}\mathbf{x}_{i}^{\top}\mathbf{w}^{*}\geq\gamma$
$=\alpha\gamma G(\mathbf{w}).$
For the third claim, by the assumption, we have
$\frac{1}{n}\cdot\frac{1}{1+\exp\big{(}y_{i}f(\mathbf{w};\mathbf{x}_{i})\big{)}}\leq G(\mathbf{w})\leq\frac{1}{2n},$
which implies that
$y_{i}f(\mathbf{w};\mathbf{x}_{i})\geq 0,\quad\forall i\in[n].$
Therefore,
$G(\mathbf{w})=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{1+\exp\bigl{(}y_{i}f(\mathbf{w};\mathbf{x}_{i})\bigr{)}}\geq\frac{1}{n}\sum_{i=1}^{n}\frac{1}{2\exp\bigl{(}y_{i}f(\mathbf{w};\mathbf{x}_{i})\bigr{)}}=\frac{1}{2}F(\mathbf{w}).$
We complete the proof of Lemma C.1. ∎
The key ingredient of the phase transition analysis is the following lemma. The main idea is to consider the gradient potential $G(\mathbf{w})$ instead of the loss function $L(\mathbf{w})$ in EoS phase. And this will decrease the order of the bound of phase transition time from $\tilde{O}(\eta^{2})$ to $\tilde{O}(\eta)$.
Lemma C.2 (A bound of $\|\mathbf{w}_{t}\|$).
For every $\eta$, we have
$\|\mathbf{w}_{t}\|\leq\sqrt{m}\cdot\frac{2+8\log(\gamma^{2}\eta t)/\alpha+8\kappa/\alpha+4\eta}{\gamma}+2\|\mathbf{w}_{0}\|.$
Proof of Lemma C.2.
By Theorem B.3, we have
$\frac{\|\mathbf{w}_{t}-\mathbf{u}\|^{2}}{2m\eta t}\leq\frac{\|\mathbf{w}_{t}-\mathbf{u}\|^{2}}{2m\eta t}+\frac{1}{t}\sum_{k=0}^{t-1}L\bigl{(}\mathbf{w}^{(k)}\bigr{)}\leq\frac{1+8\log^{2}(\gamma^{2}\eta t)/\alpha^{2}+8\kappa^{2}/\alpha^{2}+\eta^{2}}{\gamma^{2}\eta t}+\frac{\|\mathbf{w}_{0}\|^{2}}{m\eta t}.$
Besides, we know that
$\|\mathbf{u}\|^{2}\leq 2\|\mathbf{u}_{1}\|^{2}+2\|\mathbf{u}_{2}\|^{2}=\frac{4m\log(\gamma^{2}\eta t)^{2}+4m\kappa^{2}}{\alpha^{2}\gamma^{2}}+\frac{m\eta^{2}}{2\gamma^{2}}.$
Combining them, we have
$\|\mathbf{w}_{t}\|^{2}\leq 2\|\mathbf{w}_{t}-\mathbf{u}\|^{2}+2\|\mathbf{u}\|^{2}\leq m\cdot\frac{2+24\log^{2}(\gamma^{2}\eta t)/\alpha^{2}+24\kappa^{2}/\alpha^{2}+3\eta^{2}}{\gamma^{2}}+2\|\mathbf{w}_{0}\|^{2}.$
Hence, we can get a bound for $\|\mathbf{w}_{t}\|$.
$\|\mathbf{w}_{t}\|\leq\sqrt{m}\cdot\frac{2+8\log(\gamma^{2}\eta t)/\alpha+8\kappa/\alpha+4\eta}{\gamma}+2\|\mathbf{w}_{0}\|.$
Now we have completed the proof of Lemma C.2. ∎
Lemma C.3 (Gradient potential bound in the EoS phase).
For every $\eta$, we have
$\frac{1}{t}\sum_{k=0}^{t-1}G(\mathbf{w}^{(k)})\leq\frac{\langle\mathbf{w}_{t},\bar{\mathbf{w}}_{*}\rangle-\langle\mathbf{w}_{0},\bar{\mathbf{w}}_{*}\rangle}{m\alpha\gamma\eta t}\leq\frac{\sqrt{m}\|\mathbf{w}_{t}\|-\langle\mathbf{w}_{0},\bar{\mathbf{w}}_{*}\rangle}{m\alpha\gamma\eta t},\quad t\geq 1.$
Additionally, we have
$\frac{1}{t}\sum_{k=0}^{t-1}G(\mathbf{w}^{(k)})\leq\frac{2+8\log\big{(}\gamma^{2}\eta t\big{)}/\alpha+8\kappa/\alpha+4\eta}{\alpha\gamma^{2}\eta t}+\frac{3\|\mathbf{w}_{0}\|}{\alpha\gamma\eta t},\quad t\geq 1.$
Proof of Lemma C.3.
This is from the perceptron argument [Novikoff, 1962]. Specifically,
$\langle\mathbf{w}_{t+1},\bar{\mathbf{w}}_{*}\rangle$ $=\langle\mathbf{w}_{t},\bar{\mathbf{w}}_{*}\rangle-m\eta\langle\nabla L(\mathbf{w}_{t}),\bar{\mathbf{w}}_{*}\rangle$
$=\langle\mathbf{w}_{t},\bar{\mathbf{w}}_{*}\rangle-\eta\sum_{i=1}^{n}\sum_{k=1}^{m}a_{k}^{2}\ell^{\prime}(y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i}))y_{i}\phi(\mathbf{x}_{i}^{\top}\mathbf{w}_{t}^{(k)})\langle\mathbf{x}_{i},\mathbf{w}_{*}\rangle$
$\geq\langle\mathbf{w}_{t},\bar{\mathbf{w}}_{*}\rangle-\eta\sum_{i=1}^{n}\sum_{k=1}^{m}a_{k}^{2}\ell^{\prime}(y_{i}f(\mathbf{w}_{t};\mathbf{x}_{i}))\alpha\gamma$
$\geq\langle\mathbf{w}_{t},\bar{\mathbf{w}}_{*}\rangle+m\alpha\gamma\eta G(\mathbf{w}_{t}).$
Telescoping the sum, we have
$\frac{1}{t}\sum_{k=0}^{t-1}G(\mathbf{w}^{(k)})$ $\leq\frac{\langle\mathbf{w}_{t},\bar{\mathbf{w}}_{*}\rangle-\langle\mathbf{w}_{0},\bar{\mathbf{w}}_{*}\rangle}{m\alpha\gamma\eta t}$
$\leq\frac{\sqrt{m}\|\mathbf{w}_{t}\|-\langle\mathbf{w}_{0},\bar{\mathbf{w}}_{*}\rangle}{m\alpha\gamma\eta t}$
$\leq\frac{2+8\log\big{(}\gamma^{2}\eta t\big{)}/\alpha+8\kappa/\alpha+4\eta}{\alpha\gamma^{2}\eta t}+\frac{3\|\mathbf{w}_{0}\|}{\sqrt{m}\alpha\gamma\eta t}.$ by Lemma C.2
We have completed the proof of Lemma C.3. ∎
Besides, we can make use of the equivalence between $G$ and $L$ to get a bound for the loss function which is independent of the initial margin at $s$.
Lemma C.4 (A risk bound in the stable phase for Two-layer NN).
Suppose that there exists a time $s$ such that
$L(\mathbf{w}_{s})\leq\min\bigg{\{}\frac{1}{\eta(4+2\tilde{\beta})},\frac{1}{2e^{\kappa+2}n}\bigg{\}}.$
Then for every $t\geq s+1$, we have
$L(\mathbf{w}_{t})\leq\frac{2}{(t-s)\alpha^{2}\gamma^{2}}.$
Proof of Lemma C.4.
By Lemma A.3 and $f(x)$ is $\frac{1}{\sqrt{m}}$ Lipschitz and $\frac{\tilde{\beta}}{m}$ smooth, we have
$L_{k+1}\leq L_{k}-m\eta(1-(2+\tilde{\beta})\eta L(\mathbf{w}_{k}))\|\nabla L_{t}\|^{2}.$
By Lemma C.1 and $L_{t}\leq\frac{1}{\eta(4+2\tilde{\beta})}$, we have
$L_{k+1}\leq L_{k}-\frac{\alpha^{2}\gamma^{2}}{2}L_{k}^{2}.$
Multiplying $\frac{1}{L_{k}^{2}}$ in both sides, we have
$\frac{\alpha^{2}\gamma^{2}}{2}\leq\frac{L_{t}-L_{k+1}}{L_{k}^{2}}\leq\frac{1}{L_{k+1}}-\frac{1}{L_{k}}.$
Taking summation for $k=s,\ldots,t-1$, we have
$\frac{1}{L_{t}}>\frac{1}{L_{t}}-\frac{1}{L_{s}}\geq\frac{(t-s)\alpha^{2}\gamma^{2}}{2}\implies L_{t}\leq\frac{2}{(t-s)\alpha^{2}\gamma^{2}}.$
This completes the proof of Lemma C.4. ∎
At last, we will use the bound for the gradient potential to get an upper bound for the phase transition time.
C.1 Proof of Theorem 4.1
Proof of Theorem 4.1.
Applying Lemma C.3, we have
$\frac{1}{\tau}\sum_{k=0}^{\tau-1}G(\mathbf{w}^{(k)})$ $\leq\frac{2+8\log\big{(}\gamma^{2}\eta\tau\big{)}/\alpha+8\kappa/\alpha+4\eta}{\alpha\gamma^{2}\eta\tau}+\frac{3\|\mathbf{w}_{0}\|}{\sqrt{m}\alpha\gamma\eta\tau}$
[…]
Let $c_{1}=4e^{\kappa+2},c_{2}=(8+4\tilde{\beta})$. Note that we have
[…]
and that the two conditions are satisfied because
$\gamma^{2}\tau$ $\coloneqq\frac{128(1+4\kappa)}{\alpha^{2}}\max\bigg{\{}c_{2}\eta,c_{1}n,e,\frac{c_{2}\eta+c_{1}n}{\eta}\log\frac{c_{2}\eta+c_{1}n}{\eta},\frac{(c_{2}\eta+c_{1}n)\|\mathbf{w}_{0}\|}{\eta\sqrt{m}}\bigg{\}}$
$\geq\max\bigg{\{}4(2+8\kappa)\frac{c_{2}\eta+c_{1}n}{\eta\alpha^{2}},128\frac{c_{2}\eta+c_{1}n}{\eta\alpha^{2}}\log\frac{c_{2}\eta+c_{1}n}{\eta},\frac{48}{\alpha^{2}}(c_{2}\eta+c_{1}n),\frac{12}{\alpha}\frac{(c_{2}\eta+c_{1}n)}{\eta}\cdot\frac{\|\mathbf{w}_{0}\|}{\sqrt{m}}\bigg{\}}.$
So there exits $s\leq\tau$ such that
$G(\mathbf{w}_{s})\leq\min\bigg{\{}\frac{1}{e^{\kappa+2}4n},\frac{1}{{\eta}(8+4\tilde{\beta})}\bigg{\}}$
Then we have $L(\mathbf{w}_{s})\leq F(\mathbf{w}_{s})\leq 2G(\mathbf{w}_{s})\leq\Big{\{}\frac{1}{e^{\kappa+2}2n},\frac{1}{{\eta}(4+2\tilde{\beta})}\Big{\}}$. We complete the proof of Theorem 4.1. ∎
C.2 Proof of Corollary 4.2
Proof of Corollary 4.2.
The main idea is to show that $\tau\leq\frac{T}{2}$. Note that by Theorem 4.1, we have
$\tau=\frac{128(1+4\kappa)}{\alpha^{2}}\max\bigg{\{}c_{2}\eta,c_{1}n,e,\frac{c_{2}\eta+c_{1}n}{\eta}\log\frac{c_{2}\eta+c_{1}n}{\eta},\frac{(c_{2}\eta+c_{1}n)}{\eta}\cdot\frac{\|\mathbf{w}_{0}\|}{\sqrt{m}}\bigg{\}},$
in which expression $c_{1}=4e^{\kappa+2}$ and $c_{2}=(8+4\tilde{\beta})$. We can verify that,
$\frac{128(1+4\kappa)}{\alpha^{2}}c_{2}\eta$ $=\frac{128(1+4\kappa)}{\alpha^{2}}c_{2}\cdot\frac{\alpha^{2}\gamma^{2}}{256(1+4\kappa)c_{2}}T=\frac{T}{2},$
$\frac{128(1+4\kappa)c_{1}n}{\alpha^{2}}\leq\frac{T}{2}.$
Furthermore, we have $n\leq\frac{\alpha^{2}\gamma^{2}T}{256(1+4\kappa)c_{1}}$. Hence,
$\frac{c_{2}\eta+c_{1}n}{\eta}=\frac{\frac{\alpha^{2}\gamma^{2}T}{256(1+4\kappa)}+c_{1}n}{\frac{\alpha^{2}\gamma^{2}T}{256(1+4\kappa)c_{2}}}\leq\frac{2\cdot\frac{\alpha^{2}\gamma^{2}T}{256(1+4\kappa)}}{\frac{\alpha^{2}\gamma^{2}T}{256(1+4\kappa)c_{2}}}\leq 2c_{2}.$
We get that:
$\frac{128(1+4\kappa)}{\alpha^{2}}\cdot\frac{c_{2}\eta+c_{1}n}{\eta}\log\frac{c_{2}\eta+c_{1}n}{\eta}$ $\leq 2\frac{128(1+4\kappa)}{\alpha^{2}}c_{2}\ln(2c_{2})\leq\frac{128(1+4\kappa)}{\alpha^{2}}4c_{2}^{2}\leq\frac{T}{2},$
$\frac{128(1+4\kappa)}{\alpha^{2}}\cdot\frac{(c_{2}\eta+c_{1}n)}{\eta}\cdot\frac{\|\mathbf{w}_{0}\|}{\sqrt{m}}$ $\leq\frac{128(1+4\kappa)}{\alpha^{2}}\cdot 2c_{2}\frac{\|\mathbf{w}_{0}\|}{\sqrt{m}}\leq\frac{T}{2}.$
Hence, we have $\tau\leq\frac{T}{2}$. Applying Theorem 4.1, we have
$L(\mathbf{w}_{T})\leq\frac{2}{\alpha^{2}\gamma^{2}\eta(T-\tau)}\leq\frac{4}{\alpha^{2}\gamma^{2}\eta T}\leq\frac{2048(1+4\kappa)c_{2}}{\alpha^{4}\gamma^{4}T^{2}}=\mathcal{O}(1/T^{2}).$
We have completed the proof of Corollary 4.2. ∎
C.3 Proof of Theorem 4.3
Proof of Theorem 4.3.
The main idea is to construct an upper bound of $\eta$ and apply the analysis in Theorem 2.2. Note that give $\mathbf{w}_{0}=0$, we have
$f(\mathbf{w}_{0};\mathbf{x}_{i})=\frac{1}{m}\sum_{k=1}^{m}a_{k}\phi(\mathbf{x}_{i}^{\top}\mathbf{w}^{(k)}_{0})=s_{a}\phi(0),$
where $s_{a}=\sum_{k=1}^{m}a_{k}/m$. Therefore,
$[\nabla L(\mathbf{w}_{0})]^{(k)}$ $=\frac{1}{2}\ell^{\prime}(s_{a}\phi(0))\cdot\frac{a_{k}}{m}\phi^{\prime}(0)\mathbf{x}_{1}+\frac{1}{2}\ell^{\prime}(s_{a}\phi(0))\cdot\frac{a_{k}}{m}\phi^{\prime}(0)\mathbf{x}_{2}$
$=\frac{a_{k}}{m}\ell^{\prime}(s_{a}\phi(0))\phi^{\prime}(0)\frac{\mathbf{x}_{1}+\mathbf{x}_{2}}{2}$
$=\frac{a_{k}}{m}\ell^{\prime}(s_{a}\phi(0))\phi^{\prime}(0)(\gamma,\frac{\sqrt{1-\gamma^{2}}}{4}).$
Let $\bar{\mathbf{x}}\coloneqq\frac{1}{m}\ell^{\prime}(s_{a}\phi(0))\phi^{\prime}(0)(\gamma,\frac{\sqrt{1-\gamma^{2}}}{4})$, we have
$\mathbf{w}_{1}^{(k)}=0-\eta\nabla[L(\mathbf{w}_{0})]^{(k)}=-\eta a_{k}\bar{\mathbf{x}}.$
Therefore,
$f(\mathbf{w}_{1};\mathbf{x}_{i})=\frac{1}{m}\sum_{k=1}^{m}a_{k}\phi(-\mathbf{x}_{i}^{\top}(\eta a_{k}\bar{\mathbf{x}})).$
We can notice that $-\mathbf{x}_{1}^{\top}\bar{\mathbf{x}}<0$ and $-\mathbf{x}_{2}^{\top}\bar{\mathbf{x}}>0$, when $\gamma\leq 0.1$. Furthermore, we have
$f(\mathbf{w}_{1};\mathbf{x}_{1})$ $=\frac{1}{m}\sum_{a_{k}=1}\phi(-\mathbf{x}_{1}^{\top}(\eta\bar{\mathbf{x}}))+\frac{1}{m}\sum_{a_{k}=-1}-\phi(\mathbf{x}_{1}^{\top}(\eta\bar{\mathbf{x}}))$
$=\frac{1}{m}\sum_{a_{k}=1}[\phi(0)-\mathbf{x}_{1}^{\top}(\eta\bar{\mathbf{x}})\phi^{\prime}(\epsilon_{1})]+\frac{1}{m}\sum_{a_{k}=-1}[-\phi(0)-\mathbf{x}_{1}^{\top}(\eta\bar{\mathbf{x}})\phi^{\prime}(\epsilon_{2})]$
$=s_{a}\phi(0)-\eta\mathbf{x}_{1}^{\top}\bar{\mathbf{x}}\frac{1}{m}[\sum_{a_{k}=1}\phi^{\prime}(\epsilon_{1})+\sum_{a_{k}=-1}\phi^{\prime}(\epsilon_{2})]$
[…]
Note that
$\frac{1}{2}\ell(s_{a}\phi(0)-\eta\mathbf{x}_{1}^{\top}\bar{\mathbf{x}}\alpha)\leq\frac{1}{2}\ell(f(\mathbf{w}_{1};\mathbf{x}_{1}))\leq L(\mathbf{w}_{1})\leq L(\mathbf{w}_{0})=\ell(s_{a}\phi(0)).$
We apply Lemma F.7 to get
$\eta\leq\frac{|s_{a}\phi(0)|+\ln 3}{\mathbf{x}_{1}^{\top}\bar{\mathbf{x}}\alpha}.$
We use $c_{3}\coloneqq\frac{|s_{a}\phi(0)|+\ln 3}{\mathbf{x}_{1}^{\top}\bar{\mathbf{x}}\alpha}$. Now we know $\eta\leq c_{3}$. Furthermore, notice that
$\|\nabla L(\mathbf{w}_{t})\|\leq L_{t}\leq L_{0}.$
We get that $\|\mathbf{w}_{t+1}-\mathbf{w}_{t}\|\leq\eta L_{0}\leq c_{3}L_{0}$. Hence,
$|f(\mathbf{w}_{t+1};\mathbf{x}_{i})-f(\mathbf{w}_{t};\mathbf{x}_{i})|\leq c_{3}L_{0}.$
Assume that $l_{b}=\min\big{\{}\frac{1}{e^{\kappa+2}4n},\frac{1}{{\eta}\left(8+4\tilde{\beta}\right)}\big{\}}$ and
$L_{s-1}\geq l_{b},\quad L_{s}\leq l_{b}.$
We know that
$l_{b}\geq\min\bigg{\{}\frac{1}{e^{\kappa+2}4n},\frac{1}{c_{3}(8+4\tilde{\beta})}\bigg{\}}\eqqcolon l_{c}.$
We want to show that there is an lower bound for $L_{s}$. Now that
$L_{s}=\frac{1}{2}\Big{[}\ell(f(\mathbf{w}_{s};\mathbf{x}_{1}))+\ell(f(\mathbf{w}_{s};\mathbf{x}_{2}))\Big{]}.$
Applying Lemma F.6, we can get that
$L_{s}\geq\exp(-c_{3}L_{0})L_{s-1}\geq\exp(-c_{3}L_{0})l_{b}.$
Recall that by Lemma A.14, we have
$L_{t}\geq\frac{1}{\frac{1}{L_{s}}+3\tilde{\eta}\rho^{2}(t-s)},\quad t\geq s.$
Combine this with $\rho=\frac{1}{\sqrt{m}},\tilde{\eta}=\eta m$ and $L_{s}\geq\exp(-c_{3}L_{0})l_{b}$ and we get
$L_{t}\geq\frac{1}{\frac{\exp(c_{3}L_{0})}{l_{b}}+3\eta(t-s)},\quad t\geq s.$
Note that when $t\leq s$, $L_{t}\geq l_{b}$. We can get a lower bound for $L_{t}$ by
$L_{t}\geq\frac{1}{\frac{\exp(c_{3}L_{0})}{l_{b}}t+3\eta t}\geq\frac{1}{\frac{\exp(c_{3}L_{0})}{l_{b}}t+3c_{3}t}\geq\frac{1}{\frac{\exp(c_{3}L_{0})}{l_{c}}t+3c_{3}t}=\frac{c_{4}}{t},$
where $c_{4}=\frac{1}{\frac{\exp(c_{3}L_{0})}{l_{c}}+3c_{3}}$ depends only on $\{a_{j}\}_{j=1}^{m},\phi(0),\kappa,\tilde{\beta}$ and $n$. Now we have completed the proof of Theorem 4.3. ∎
Appendix D Scaling and Homogenous Error
In this section, we consider different scaling of two-layer networks. We add a scaling factor $b$ into the model, i.e.,
$f(\mathbf{w};\mathbf{x})=\frac{b}{m}\sum_{j=1}^{m}a_{j}\phi(\mathbf{x}^{\top}\mathbf{w}^{(j)}).$
We will show that given a limited computation budget $T$ (total iterations), larger $b$ and a corresponding best $\tilde{\eta}=\eta\cdot m$ will achieve the same best rate as $b=1$, i.e., $O(1/T^{2})$. While for smaller $b$, the rate is $O(b^{-3}/T^{2})$. Before we present the analysis, here are the bounds with $b$ and $\tilde{\eta}=m\cdot\eta$ following the process of Lemma C.3:
$\frac{1}{t}\sum_{k=0}^{t-1}L(\mathbf{w}_{k})\leq\frac{1+8\log^{2}(\gamma^{2}\eta t)/(\alpha^{2}b^{2})+8\kappa^{2}/\alpha^{2}+\eta^{2}b^{2}}{\gamma^{2}\eta t}+\frac{\|\mathbf{w}_{0}\|^{2}}{m\eta t},$
$\frac{1}{t}\sum_{k=0}^{t-1}G(\mathbf{w}_{k})\leq\frac{2+8\log(\gamma^{2}\eta t)/(\alpha b)+8\kappa/\alpha+2\eta b}{\alpha\gamma^{2}b\eta t}+\frac{3\|\mathbf{w}_{0}\|}{\sqrt{m}\eta bt}.$
Case when $b\geq 1$.
Given the previous bounds, we have the following results following the idea in Appendix C:
  • Gradient potential bound: $G(\mathbf{w}_{t})\leq\frac{C}{t}$ for all $t\geq 0$,
  • Phase transition threshold: $G(\mathbf{w}_{s})\leq\min\Big{\{}1/4e^{\kappa+2}n,1/\eta(8\rho^{2}b^{2}+4\tilde{\beta}b)\Big{\}}$,
  • Stable phase bound: $L(\mathbf{w}_{t})\leq\frac{2}{Cb^{2}\eta(t-s)}$,
where $C$ depends on $\alpha,\gamma$. Combine the first two arguments and assume $\eta(8\rho^{2}b^{2}+4\tilde{\beta}b)\geq 4e^{\kappa+2}n$. We get $s\leq C\eta(8\rho^{2}b^{2}+4\tilde{\beta}b)$. Plug this into the third bound. We have
$L(\mathbf{w}_{T})\leq\frac{2}{Cb^{2}\eta(T-C\eta(8\rho^{2}b^{2}+4\tilde{\beta}b))}.$
It’s obvious that the best $\eta=\frac{T}{16\rho^{2}b^{2}C+8\tilde{\beta}bC}$. Hence,
$L(\mathbf{w}_{T})\leq\frac{8(8\rho^{2}b^{2}C+4\tilde{\beta}bC)}{Cb^{2}T^{2}}=\mathcal{O}\bigg{(}\frac{1}{T^{2}}\bigg{)}.$
Then, the rate is still $O(1/T^{2})$.
Case when $b<1$.
Similarly, we can get the following bounds:
  • Gradient potential bound: $G(\mathbf{w}_{t})\leq\frac{Cb^{-2}}{t}$ for all $t\geq 0$,
  • Phase transition threshold: $G(\mathbf{w}_{s})\leq\min\Big{\{}1/4e^{\kappa+2}n,1/\eta(8\rho^{2}b^{2}+4\tilde{\beta}b)\Big{\}}$,
  • Stable phase bound: $L(\mathbf{w}_{t})\leq\frac{2}{Cb^{2}\eta(t-s)}$,
where $C$ depends on $\alpha,\gamma$. Without loss of generality, we can assume $\eta(8\rho^{2}b^{2}+4\tilde{\beta}b)\geq 4e^{\kappa+2}n$, since $\eta$ can be small enough. Then, we have
$s\leq C\eta(8\rho^{2}+4\tilde{\beta}b^{-1}).$
Then, we can get
$L(\mathbf{w}_{T})\leq\frac{2}{Cb^{2}\eta(T-C\eta(8\rho^{2}+4\tilde{\beta}b^{-1}))}.$
It’s obvious that the best $\eta=\frac{T}{16\rho^{2}C+8\tilde{\beta}b^{-1}C}$. Hence,
$L(\mathbf{w}_{T})\leq\frac{8(8\rho^{2}Cb^{-2}+4\tilde{\beta}b^{-3}C)}{CT^{2}}=\mathcal{O}\bigg{(}\frac{b^{-3}}{T^{2}}\bigg{)}.$
Combining the analysis for two cases, we observe that when $b\geq 1$, the fast loss rate is $\mathcal{O}(1/T^{2})$ given finite budget $T$. While $b<1$, the rate is $\mathcal{O}(b^{-3}/T^{2})$. In our main results, we set $b=1$ for the mean-field scaling. Under the mean-field regime, all bounds are independent of the number of neurons since we consider the dynamics of the distributions of neurons. Alternatively, if we set $b=\sqrt{m}$, then the model becomes:
$f(\mathbf{w};\mathbf{x})=\frac{1}{\sqrt{m}}\sum_{j=1}^{m}a_{j}\phi(\mathbf{x}^{\top}\mathbf{w}^{(j)}).$
The model falls into the NTK regime. The loss threshold will be related to $m$, but the loss rate is the same as that of the mean-field scaling.
Appendix E Additional Proofs
E.1 Proof of Example 2.1
Proof of Example 2.1.
Recall that the two-layer neural network is defined as:
$f(\mathbf{w};\mathbf{x})=\frac{1}{m}\sum_{j=1}^{m}a_{j}\phi(\mathbf{x}^{T}\mathbf{w}^{(j)}).$
We can verify that if $\phi(x)$ is $\beta$-smooth and $\rho$-Lipschitz with respect to $x$, then $f(\mathbf{w};\mathbf{x})$ is $\beta/m$-smooth and $\rho/\sqrt{m}$-Lipschitz with respect to $\mathbf{w}$. This is because:
$\nabla L(\mathbf{w})$ $=\hat{\mathbb{E}}\ell^{\prime}(yf(\mathbf{w};\mathbf{x}))y\nabla f(\mathbf{w};\mathbf{x}),$
$\nabla^{2}L(\mathbf{w})$ $=\hat{\mathbb{E}}\ell^{\prime\prime}(yf(\mathbf{w};\mathbf{x}))\nabla f(\mathbf{w};\mathbf{x})^{\otimes 2}+\ell^{\prime}(yf(\mathbf{w};\mathbf{x}))y\nabla^{2}f(\mathbf{w};\mathbf{x}),$
and that
$\nabla f(\mathbf{w};\mathbf{x})=\begin{pmatrix}\vdots\\ \frac{1}{m}a_{j}\phi^{\prime}(\mathbf{x}^{\top}\mathbf{w}^{(j)})\mathbf{x}\\ \vdots\end{pmatrix},\quad\nabla^{2}f(\mathbf{w};\mathbf{x})=\begin{pmatrix}\ddots&0&0\\ 0&\frac{1}{m}a_{j}\phi^{\prime\prime}(\mathbf{x}^{\top}\mathbf{w}^{(j)})\mathbf{x}\mathbf{x}^{\top}&0\\ 0&0&\ddots\end{pmatrix}.$
Now, we will focus on the parameters of each activation function.
  • GELU. $\phi(x)=x\cdot\mathrm{erf}(1+(x/\sqrt{2}))/2=x\cdot F(x)$.
    $\phi^{\prime}(x)$ $=F(x)+x\cdot f(x),$
    $\phi^{\prime\prime}(x)$ $=2f(x)+x\cdot f^{\prime}(x),$
    where $F(x),f(x)$ are the CDF and PDF of standard normal distribution. Note that $xf(x)=\frac{x}{\sqrt{2\pi}}e^{-x^{2}/2}$ and $(xf(x))^{\prime}=\frac{1}{\sqrt{2\pi}}(1-x^{2})e^{-x^{2}/2}$. We can find the maximum of $xf(x)$ is $\frac{1}{\sqrt{2\pi}}e^{-1/2}$. Besides, we know that $F(x),f(x)\leq 1$ and $x\cdot f^{\prime}(x)\leq 0$. Combining them, we have $\rho=1+e^{-1/2}/\sqrt{2\pi}$ and $\beta=2$. For $\kappa$, $\phi-\phi^{\prime}(x)x=-x\cdot f(x)$. So the bound of $\kappa$ is $e^{-1/2}/\sqrt{2\pi}$.
  • Softplus. $\phi(x)=\log(1+e^{x})$. Therefore,
    $\phi^{\prime}(x)=\frac{e^{x}}{1+e^{x}}\leq 1,$
    $\phi^{\prime\prime}(x)=\frac{e^{x}}{(1+e^{x})^{2}}\leq 1.$
    Besides,
    $(\phi(x)-\phi^{\prime}(x)x)^{\prime}=\bigg{(}\log(1+e^{x})-\frac{e^{x}x}{1+e^{x}}\bigg{)}^{\prime}=-\frac{e^{x}x}{(1+e^{x})^{2}}.$
    So the maximum is $\phi(0)-\phi^{\prime}(0)0=\log 2$. Besides, when $x>1$, $\phi(x)\geq x\geq\frac{e^{x}x}{1+e^{x}}$. When $x\to-\infty$, $\phi(x)-\phi^{\prime}(x)x\to 0$. Therefore, $\kappa=\log 2$.
  • Sigmoid. $\phi(x)=1/(1+e^{-x})$. Hence,
    $\phi^{\prime}(x)=\frac{e^{-x}}{(1+e^{-x})^{2}}\leq 1,$
    $\phi^{\prime\prime}(x)=\frac{e^{-2x}-e^{-x}}{(1+e^{-x})^{3}}\leq 1.$
    As for $\kappa$, we know that
    $|\phi(x)-\phi^{\prime}(x)x|=\bigg{|}\frac{1+e^{-x}-xe^{-x}}{(1+e^{-x})^{2}}\bigg{|}\leq\frac{1+e^{-x}+|x|e^{-x}}{(1+e^{-x})^{2}}.$
    Note that $|x|e^{-x}\leq e^{-2x}+1$. We have
    $|\phi(x)\leq\frac{1+e^{-x}+e^{-2x}+1}{(1+e^{-x})^{2}}\leq 2.$
  • Tanh. $\phi(x)=\frac{e^{x}-e^{-x}}{e^{x}+e^{-x}}\leq 1$. Note that
    $\phi^{\prime}(x)=1-\phi(x)^{2}\leq 1$
    $\phi^{\prime\prime}(x)=2\phi(x)^{3}-\phi(x)\leq 2.$
    Besides, we know that
    $|x\phi^{\prime}(x)|=\frac{4|x|}{(e^{x}+e^{-x})^{2}}\leq 4.$
    Hence
    $|\phi(x)-\phi^{\prime}(x)|\leq|\phi(x)|+|x\phi^{\prime}(x)|\leq 5.$
  • SiLU. Note that
    $\phi^{\prime}(x)=\frac{1+e^{-x}+xe^{-x}}{(1+e^{-x})^{2}},$
    $\phi^{\prime\prime}(x)=\frac{(2-x)e^{-x}}{(1+e^{-x})^{2}}+\frac{xe^{-2x}}{(1+e^{-x})^{3}}.$
    Because $|x|e^{-x}\leq e^{-2x}+1$ and $|x|e^{-2x}\leq e^{-3x}+1$. We get $|\phi^{\prime}(x)|\leq 2$ and $\phi^{\prime\prime}(x)|\leq 4$. At last,
    $|\phi(x)-x\phi^{\prime}(x)|=\frac{|x|e^{-x}}{(1+e^{-x})^{2}}\leq 1.$
  • Huberized ReLU. It’s obvious that $\phi^{\prime}(x)\leq 1$ and $\beta=1/h$. Note that $\phi$ is not second-order differentiable. At last,
    $|\phi(x)-x\phi^{\prime}(x)|=\begin{cases}0&x<0,\\ x^{2}/2h&0\leq x\leq h,\\ h/2&x>h.\end{cases}$
    Hence, it’s upper bounded by $h/2$. Now we have completed the proof of Example 2.1.
E.2 Proof of Example 3.1
Proof of Example 3.1.
Because for activation functions in Example 2.1, $\beta\leq 4$ and $\rho\leq 2$. Hence, for $\tilde{\phi}(x)=cx+(1-c)\phi(x)/4$, $\tilde{\beta}=1$ and $\rho=1$. Besides, since $10.5<c<$ 1, we must have $(\tilde{\phi}(x))^{\prime}\geq 0.25$. ∎
Appendix F Additional Lemmas
Lemma F.1.
If $\frac{1}{2}\geq L_{1}\geq\frac{1}{c}$ and $L_{2}\geq L_{1}-L_{1}^{2}$, we have
$L_{2}\geq\frac{1}{c+2}.$
Proof of Lemma F.1.
For function $g(x)=x-x^{2}$, $g^{\prime}(x)=1-2x$. If $x\leq\frac{1}{2}$, then $g(x)$ is increasing. Then
$g(L_{1})\geq g(\frac{1}{c})=\frac{c-1}{c^{2}}=\frac{c^{2}+c-2}{c^{2}(c+2)}\geq\frac{1}{c+2}.$
Now we have completed the proof of Lemma F.1. ∎
Lemma F.2.
Given a continuous function $f(x)$ s.t. $|f(x)-\langle\nabla f(x),x\rangle|\leq\kappa$, then for a fixed constant $r>0$ there exists $C_{r,\kappa}$ and $C_{r}$ s.t. for any $\|x\|\geq r$,
$|f(x)|\leq C_{r,\kappa}\|x\|,$
and for any $x$,
$|f(x)|\leq C_{r,\kappa}\|x\|+C_{r}.$
Proof of Lemma F.2.
Since $f$ is continuous, let
$C_{r}=\max_{\|x\|=r}|f(x)|/r.$
Now for any $\|x\|>r$, let $y=\frac{rx}{\|x\|}$ and consider $g(s)=\frac{f(sy)}{s}$. Then we have
$g^{\prime}(s)=\frac{\langle\nabla f(sy),sy\rangle-f(sy)}{s^{2}}.$
Therefore, $-\frac{\kappa}{s^{2}}\leq g^{\prime}(s)\leq\frac{\kappa}{s^{2}}$. Let $s=\|x\|/r$,
[…]
$\leq rC_{r}+\kappa.$
Therefore, $f(x)\leq(C_{r}+\frac{\kappa}{r})\cdot\|x\|.$ Similarly, we can show that $-f(x)\leq(C_{r}+\frac{\kappa}{r})\cdot\|x\|$. Therefore, for any $\|x\|\geq r$,
$|f(x)|\leq(C_{r}+\frac{\kappa}{r})\cdot\|x\|.$
Let $D=\max_{\|x\|\leq r}|f(x)|$, we have for any $x$,
$|f(x)|\leq(C_{r}+\frac{\kappa}{r})\cdot\|x\|+D.$
We have completed the proof of Lemma F.2. ∎
Lemma F.3.
Fixing $c>1$, then for every $0<x\leq\frac{1}{c}$, we have
$x\leq\frac{-1}{\log(cx)}.$
Proof of Lemma F.3.
This is equivalent to show that
$x\log(cx)\geq-1.$
Let $s(x)=x\log(cx)$, then $s^{\prime}(x)=1+\log(cx)$. Hence $s(x)$ is decreasing when $0<x<\frac{1}{ce}$ and is increasing when $x\geq\frac{1}{ce}$. The minimum of $s(x)$ is achieved at $x=\frac{1}{ce}$, which is
$s(1/(ce))=-\frac{1}{ce}\geq-1.$
This completes the proof of Lemma F.3. ∎
Lemma F.4.
Given $0<b\leq\frac{1}{2}$ and $a>0$, we have
$\frac{1+a}{1-b}\leq(1+2a+2b).$
Proof of Lemma F.4.
This is equivalent to show that
$(1+a)\leq(1+2a+2b)(1-b)=1+2a+b-2ab-2b^{2}.$
This is equivalent to
$2b(a+b)\leq(a+b).$
Since $a+b>0$ and $b\leq\frac{1}{2}$, this is true. Now we have completed the proof of Lemma F.4. ∎
Lemma F.5.
Given $c>e$, we have for any $x>2c\log c$,
$\frac{\log x}{x}\leq\frac{1}{c}.$
Proof of Lemma F.5.
It’s equivalent to show that $x-c\log x\geq 0$. Let $g(x)=x-c\log x$. $g^{\prime}(x)=1-c/x$. When $x>2c\log c>2c$, $g^{\prime}(x)<0$. Hence, the minimal is $g(2c\log c)$. Note that
$g(2c\log c)=2c\log c-c\log c-c\log 2-c\log\log c=c\log c-c\log 2-c\log\log c=c\log\frac{c}{2\log c}.$
Now we want to show that $c>2\log c$. Let $h(y)=y-2\log y$. $h^{\prime}(y)=1-2/y>0$ when $y>e$. $h(e)=e-2>0$. Hence $h(c)>h(e)>0$ and $g(2c\log c)>0$. This leads to $g(x)>0.$ Then, we complete the proof of Lemma F.5. ∎
Lemma F.6.
Given $\ell(x)=\log(1+e^{-x})$ and $c>0$, we have for any $x$,
$\ell(x+c)\geq\exp(-c)\ell(x).$
Proof of Lemma F.6.
Let $g(x)=\ell(x+c)-\exp(-c)\ell(x)$. Then, we have
$g^{\prime}(x)=\frac{-1}{1+\exp(x+c)}+\frac{1}{\exp(c)+\exp(x+c)}<0.$
Therefore, $g(x)$ is monotonically decreasing. When $x\to\infty$, we have
$\lim_{x\to\infty}g(x)=\lim_{x\to\infty}[\ell(x+c)-\exp(-c)\ell(x)]=\exp(-x-c)-\exp(-c)\exp(-x)=0.$
Therefore, $g(x)\geq 0$ for any $x$. Now, we complete the proof of Lemma F.6. ∎
Lemma F.7.
Assume $\ell(x)=\log(1+e^{-x})$. If $\ell(x+c)\leq 2\ell(x)$, we have
$c\leq\ln 3+|x|.$
Proof of Lemma F.7.
Note that
$\ell(x+c)-2\ell(x)=\log\frac{1+e^{x+c}}{1+2e^{x}+e^{2x}}\leq 0.$
Then,
$\frac{1+e^{x+c}}{1+2e^{x}+e^{2x}}\leq 1\implies e^{c}\leq 2+e^{x}\leq 2+e^{|x|}\leq 3e^{|x|}.$
Therefore, $c\leq\ln 3+|x|$. This completes the proof of Lemma F.7. ∎